### v1

# Two-grid convergence theory for symmetric positive semidefinite linear systems

Abstract.
This paper is devoted to the convergence theory of two-grid methods for symmetric positive semidefinite linear systems, with particular focus on the singular case. In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for characterizing the convergence factor of two-grid methods. More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers, including both linear and general cases. A key feature of our analysis is that it does not require any additional assumptions on the system matrix, especially on its null space.
Key words and phrases:
two-grid methods, symmetric positive semidefinite matrices, singular systems, convergence analysis
2020 Mathematics Subject Classification:
Primary 65F08, 65F10, 65N55; Secondary 15A10, 15A18
1. Introduction
As a powerful solver with linear or near-linear computational complexity, multigrid has been widely used for the numerical solution of partial differential equations; see, e.g., [11, 4, 26, 27, 31]. The foundation of multigrid is a two-grid scheme, which consists of two complementary error-reduction processes: smoothing (or relaxation) and coarse-grid correction. In the correction step, a coarse-grid system with fewer unknowns needs to be solved. In practice, it is often too costly to solve the coarse-grid system exactly. A typical strategy is to apply the two-grid scheme recursively in the correction steps, which will yield an algorithm with multilevel structure (i.e., multigrid method). For symmetric positive definite (SPD) linear systems, the convergence properties of two-grid and multigrid methods have been well studied; see, e.g., [29, 35, 30, 7, 8, 17, 31, 32, 33, 34].
Many systems arising in mathematical models may only be semidefinite, such as discretized partial differential equations with Neumann boundary conditions (see, e.g., [1]) and Laplacian matrices of graphs (see, e.g., [24, 16]). In such cases, the system matrix, denoted by $A$, is singular and symmetric positive semidefinite (SPSD). Some convergence properties of classical iterative methods for SPSD systems can be found, e.g., in [13, 5, 14, 6, 9, 28, 10]. The convergence of Schwarz methods for SPSD systems was studied in [21]. Our focus here is on the convergence of two-grid methods for SPSD systems. In practice, two-grid convergence theory can be used to guide or motivate the design of multigrid methods; see, e.g., [3, 19, 18]. Moreover, two-grid convergence estimates can be applied to the analysis of multigrid methods; see, e.g., [33, 34].
In the SPD case, an identity for the convergence factor of two-grid methods was established in [8]; see [30, 36] for an abstract version. Ideally, one can derive a two-grid convergence identity for the semidefinite case without imposing any restrictions on the null space of $A$, which is denoted by $\mathcal{N}(A)$. Under several essential assumptions on local null spaces, a sharp convergence estimate for the successive subspace correction methods [29, 30] applied to SPSD systems was presented in [15]; see [28] for a simplified analysis. More specifically, [15] offers an identity for the seminorm of error propagation operator, despite it involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement. In [2], a classical two-grid convergence estimate for SPD systems [20, 23, 25] was extended to the semidefinite case. A two-grid convergence theory for singular systems was developed in [22], in which a restrictive condition on the coarse space and $\mathcal{N}(A)$ is required. Even in the special SPSD case, the two-grid estimates in [22] involve an extended prolongation operator, whose construction is based on $\mathcal{N}(A)$. A noteworthy contribution is the two-grid convergence identity appearing in [31, Theorem 5.3], which was established under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
The aforementioned two-grid convergence theory and the associated assumptions rely largely on $\mathcal{N}(A)$, while the properties of $\mathcal{N}(A)$ may not be easily available. Consequently, it might be tough to apply those results or check the required conditions. Motivated by this observation, we establish a new convergence theory for two-grid methods applied to SPSD systems, which lifts the restrictions on $\mathcal{N}(A)$. Our main contribution is twofold.
  • Firstly, we derive a succinct identity for the convergence factor of two-grid methods, provided that the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver (it is called the exact coarse solver).
  • Secondly, we present some convergence estimates for two-grid methods with inexact coarse solvers, including both linear and general cases.
It is worth mentioning that our theory has several distinctive features in comparison with the existing ones, as described below.
  • It does not require any additional assumptions on $A$, especially on $\mathcal{N}(A)$.
  • It yields an identity for the convergence factor of exact two-grid methods instead of an upper bound.
  • It offers some convergence bounds for two-grid methods with inexact coarse solvers.
The rest of this paper is organized as follows. In Section 2, we give a sufficient and necessary condition for two-grid convergence factor to be less than $1$. In Section 3, we present a concise identity for the convergence factor of exact two-grid methods. In Section 4, we study the convergence of two-grid methods with approximate coarse solvers. In Section 5, we make some concluding remarks.
2. Preliminaries
We start with some notation that will be used in the subsequent discussions.
  • For two symmetric matrices $S_{1}$ and $S_{2}$, the symbol $S_{1}\succ S_{2}$ (resp., $S_{1}\succeq S_{2}$) means that $S_{1}-S_{2}$ is positive definite (resp., positive semidefinite).
  • $I_{n}$ denotes the $n\times n$ identity matrix (or $I$ when its size is clear from context).
  • $\mathcal{R}(\cdot)$ denotes the range (or column space) of a matrix.
  • $\mathcal{N}(\cdot)$ denotes the null space of a matrix.
  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix, respectively.
  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix.
  • $\lambda(\cdot)$ denotes the spectrum of a matrix.
  • $\|\cdot\|_{2}$ denotes the Euclidean norm of a vector.
  • $(\cdot)^{\dagger}$ denotes the Moore–Penrose inverse of a matrix.
For any $A\in\mathbb{R}^{n\times n}\backslash\{0\}$ and $A\succeq 0$, we define
$\|\mathbf{v}\|_{A}:=\sqrt{\mathbf{v}^{T}A\mathbf{v}}\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
(2.1) $\|X\|_{A}:=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\|X\mathbf{v}\|_{A}}{\|\mathbf{v}\|_{A}}\quad\forall\,X\in\mathbb{R}^{n\times n}.$
It is easy to check that the function $\|\cdot\|_{A}:\mathbb{R}^{n\times n}\rightarrow[0,+\infty]$ satisfies the following properties:
  • (a)
    $\|X\|_{A}\geq 0$ for all $X\in\mathbb{R}^{n\times n}$;
  • (b)
    $\|\alpha X\|_{A}=|\alpha|\|X\|_{A}$ for all $\alpha\in\mathbb{R}$ and $X\in\mathbb{R}^{n\times n}$;
  • (c)
    $\|X_{1}+X_{2}\|_{A}\leq\|X_{1}\|_{A}+\|X_{2}\|_{A}$ for all $X_{1},X_{2}\in\mathbb{R}^{n\times n}$.
That is, (2.1) defines a seminorm on $\mathbb{R}^{n\times n}$.
The following lemma gives two simple but useful properties of the $A$-seminorm.
Lemma 2.1.
If $X\in\mathbb{R}^{n\times n}$ and $A^{\frac{1}{2}}X=YA^{\frac{1}{2}}$ for some $Y\in\mathbb{R}^{n\times n}$, then
(2.2) $\|X\mathbf{v}\|_{A}\leq\|X\|_{A}\|\mathbf{v}\|_{A}\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
(2.3) $\|X\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
Proof.
Note that
$\mathcal{N}(A)=\mathcal{N}\big{(}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad\mathcal{R}(A)=\mathcal{R}\big{(}A^{\frac{1}{2}}\big{)}.$
It can be readily seen from (2.1) that (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
$\|X\mathbf{v}\|_{A}=\big{\|}A^{\frac{1}{2}}X\mathbf{v}\big{\|}_{2}=\big{\|}YA^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}=0,$
where we have used the condition $A^{\frac{1}{2}}X=YA^{\frac{1}{2}}$. Hence, the inequality (2.2) holds for all $\mathbf{v}\in\mathbb{R}^{n}$.
From (2.1), we have
$\|X\|_{A}$ $=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big{\|}A^{\frac{1}{2}}X\mathbf{v}\big{\|}_{2}}{\big{\|}A^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big{\|}YA^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}{\big{\|}A^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}$
$=\sup_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}},$
which yields the expression (2.3). ∎
Consider the solution of linear systems of the form
(2.4) $A\mathbf{u}=\mathbf{f},$
where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$, $A\succeq 0$, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$. When $A$ is singular (namely, $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$. We aim to solve (2.4) using a two-grid method, the output of which is an approximation to a solution to (2.4). Two very basic assumptions involved in the two-grid analysis are made below.
  • Let $M$ be an $n\times n$ matrix such that $\|I-MA\|_{A}\leq 1$.
  • Let $P$ be an $n\times n_{\rm c}$ prolongation (or interpolation) matrix such that the Galerkin coarse-grid matrix $A_{\rm c}:=P^{T}AP$ is nonzero, where $n_{\rm c}\,(<n)$ is the number of coarse variables.
Remark 2.2.
We remark that $M$ is not necessarily nonsingular. Moreover, measured in $\|\cdot\|_{A}$, the smoothing iteration matrix in Algorithm 1 below (i.e., $I-MA$) is not required to be a strict contraction. Due to
$A^{\frac{1}{2}}(I-MA)=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from (2.3) that
$\|I-MA\|_{A}^{2}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}^{2}}{\|\mathbf{v}\|_{2}^{2}}=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(2.5) $\overline{M}:=M+M^{T}-M^{T}AM.$
Thus, $\|I-MA\|_{A}\leq 1$ if and only if $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$, or, equivalently, $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathbb{R}^{n}\backslash\{0\}$ (i.e., $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$), because $\mathbb{R}^{n}$ is the direct sum of $\mathcal{R}(A)$ and $\mathcal{N}(A)$, denoted by $\mathbb{R}^{n}=\mathcal{R}(A)\oplus\mathcal{N}(A)$.
With the above assumptions, a two-grid method for solving (2.4) is described by Algorithm 1, in which the Moore–Penrose inverse $A_{\rm c}^{\dagger}$ is used as a coarse solver. For convenience, such an algorithm is called an exact two-grid method.
Algorithm 1  Exact two-grid method.
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$triangleright $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big{(}\mathbf{f}-A\mathbf{u}^{(1)}\big{)}$
3:Coarse-grid correction: $\mathbf{e}_{\rm c}\leftarrow A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
4:Prolongation: $\mathbf{u}_{\rm TG}\leftarrow\mathbf{u}^{(1)}+P\mathbf{e}_{\rm c}$
Remark 2.3.
The output of Algorithm 1 converges to a particular solution to (2.4) as $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}$ tends to zero. Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}=A\mathbf{u}=\mathbf{f}$.
From Algorithm 1, we have
(2.6) $\mathbf{u}-\mathbf{u}_{\rm TG}=E_{\rm TG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)},$
where
$E_{\rm TG}=\big{(}I-PA_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA).$
Define
(2.7) $\Pi:=A^{\frac{1}{2}}PA_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}.$
It can be easily verified that
$\Pi^{T}=\Pi=\Pi^{2},\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)},\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}.$
Thus, $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}$ onto $\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}$. Since
$A^{\frac{1}{2}}E_{\rm TG}=(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from Lemma 2.1 and (2.6) that
$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
where $\|E_{\rm TG}\|_{A}$, called the convergence factor of Algorithm 1, is given by
(2.8) $\|E_{\rm TG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm TG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(2.9) $F_{\rm TG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Hence, it holds that $\|E_{\rm TG}\|_{A}<1$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
The following lemma provides a sufficient and necessary condition (namely, (2.10)) for $\|E_{\rm TG}\|_{A}<1$. It is worth pointing out that (2.10) is not really a restriction on $\mathcal{N}(A)$. In view of (2.10), a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$, which in fact can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.6. Throughout this paper, we always assume that (2.10) holds.
Lemma 2.4.
Under the assumptions of Algorithm 1, $\|E_{\rm TG}\|_{A}<1$ if and only if
(2.10) $\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
Proof.
Observe first that
$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq 0}+\underbrace{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}_{\geq 0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
$\mathcal{N}(A)\subseteq\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}.$
[…]
$0=P^{T}(I-AM)A^{\frac{1}{2}}\mathbf{x}=P^{T}(I-AM)A^{\frac{1}{2}}\mathbf{x}_{1}=P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1},$
it follows that
$\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1}=0.$
In addition,
$A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}_{1}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}(\mathbf{x}_{1}+\mathbf{x}_{2})=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}=0.$
Hence,
$\mathbf{x}_{1}^{T}F_{\rm TG}\mathbf{x}_{1}=\mathbf{x}_{1}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}_{1}+\mathbf{x}_{1}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1}=0,$
which contradicts with the condition $\|E_{\rm TG}\|_{A}<1$, or, equivalently, $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
(2.10)$\Rightarrow\|E_{\rm TG}\|_{A}<1$”: Conversely, if (2.10) holds, then $\|E_{\rm TG}\|_{A}<1$. Otherwise, there must exist a vector $\mathbf{y}\in\mathcal{R}(A)\backslash\{0\}$ such that $\mathbf{y}^{T}F_{\rm TG}\mathbf{y}=0$. Then
$\mathbf{y}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{y}=0\quad\text{and}\quad\mathbf{y}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{y}=0,$
which lead to
$\mathbf{y}\in\mathcal{N}\big{(}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
and
$\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{y}\in\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
Thus,
$\mathbf{y}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)},$
which contradicts with (2.10), because $\mathbf{y}\notin\mathcal{N}(A)$. This completes the proof. ∎
Corollary 2.5.
Under the assumptions of Algorithm 1, if $\overline{M}\succeq 0$ and
(2.11) $\mathcal{N}(\overline{M})\cap\mathcal{R}(A)=\{0\},$
then $\|E_{\rm TG}\|_{A}<1$.
Proof.
For any $\mathbf{v}\in\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}$, we have
$A^{\frac{1}{2}}\mathbf{v}\in\mathcal{N}(\overline{M})\cap\mathcal{R}(A).$
It then follows from (2.11) that $\mathbf{v}\in\mathcal{N}(A)$. Since $\mathbf{v}$ is arbitrary and
$\mathcal{N}(A)\subseteq\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)},$
we obtain
$\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
which, together with the positive semidefiniteness of $\overline{M}$, yields
$\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\overline{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
The desired result then follows from Lemma 2.4. ∎
Remark 2.6.
From Corollary 2.5, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, provided that the diagonal entries of $A$ are all nonzero. Let
$A=D+L+L^{T},$
where $D$ and $L$ denote the diagonal and strictly lower triangular parts of $A$, respectively. Without loss of generality, we assume that $D\succ 0$. If some diagonal entry of $D$ is zero, the corresponding row and column in $A$ are both zero, in which case one may solve a reduced version of (2.4) obtained by deleting the zero row and column. If $M=\omega D^{-1}$ with $0<\omega<\frac{2}{\lambda_{\max}(D^{-1}A)}$, then
$\overline{M}=\omega^{2}D^{-1}\big{(}2\omega^{-1}D-A\big{)}D^{-1}\succ 0.$
Another typical choice is $M=(D+L)^{-1}$, in which case one has
$\overline{M}=(D+L)^{-T}D(D+L)^{-1}\succ 0.$
3. Convergence analysis of Algorithm 1
In this section, we present a succinct identity for characterizing the convergence factor $\|E_{\rm TG}\|_{A}$.
Define
$r:=\operatorname*{rank}(A)\quad\text{and}\quad s:=\operatorname*{rank}(A_{\rm c}).$
Since $A_{\rm c}=P^{T}AP$, it follows that $s\leq r$. In the extreme case $s=r$, we have
$\operatorname*{rank}(\Pi)=\operatorname*{rank}\big{(}A^{\frac{1}{2}}P\big{)}=s=r,$
which, combined with the fact $\mathcal{R}(\Pi)\subseteq\mathcal{R}(A)$, leads to
$\mathcal{R}(\Pi)=\mathcal{R}(A).$
Then
$(I-\Pi)\mathbf{v}=0\quad\forall\,\mathbf{v}\in\mathcal{R}(A),$
which, together with (2.8), yields $\|E_{\rm TG}\|_{A}=0$. Such an extreme case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
We first prove a technical lemma, which gives a reformulation of the convergence factor $\|E_{\rm TG}\|_{A}$.
Lemma 3.1.
Under the assumptions of Algorithm 1 and the condition (2.10), one has
(3.1) $\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm TG})},$
where $F_{\rm TG}$ is given by (2.9).
Proof.
It can be easily seen from (2.7) and (2.9) that
(3.2) $\mathcal{N}(A)\subseteq\mathcal{N}(F_{\rm TG}).$
For any $\mathbf{v}\in\mathcal{N}(F_{\rm TG})$, we have
$0=\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}+\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v},$
from which we deduce that
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\in\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
Then
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
where we have used the condition (2.10). The arbitrariness of $\mathbf{v}$ implies that
(3.3) $\mathcal{N}(F_{\rm TG})\subseteq\mathcal{N}(A).$
Combining (3.2) and (3.3), we obtain
(3.4) $\mathcal{N}(F_{\rm TG})=\mathcal{N}(A)$
and hence
$\mathcal{R}(F_{\rm TG})=\mathcal{R}(A).$
Since $F_{\rm TG}\succeq 0$ and $\operatorname*{rank}(F_{\rm TG})=r$, there exists an orthogonal matrix $Q_{1}\in\mathbb{R}^{n\times n}$ such that
$Q_{1}^{T}F_{\rm TG}Q_{1}=\begin{pmatrix}\Lambda_{r}&0\\ 0&0\end{pmatrix},$
where $\Lambda_{r}\in\mathbb{R}^{r\times r}$ is diagonal and $\Lambda_{r}\succ 0$. We then have
$\mathcal{R}(A)=\mathcal{R}(F_{\rm TG})=\mathcal{R}\bigg{(}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\bigg{)}.$
Thus,
$\|E_{\rm TG}\|_{A}^{2}$ $=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
[…]
$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\Lambda_{r}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}$
$=1-\lambda_{\min}(\Lambda_{r})$
$=1-\lambda_{n-r+1}(F_{\rm TG}),$
which yields the expression (3.1). ∎
Remark 3.2.
A symmetric two-grid scheme can be obtained by adding the postsmoothing step
$\mathbf{u}_{\rm STG}\leftarrow\mathbf{u}_{\rm TG}+M^{T}(\mathbf{f}-A\mathbf{u}_{\rm TG})$
after the prolongation step of Algorithm 1. The error propagation of the symmetric two-grid scheme can be expressed as
$\mathbf{u}-\mathbf{u}_{\rm STG}=E_{\rm STG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
with
$E_{\rm STG}=(I-M^{T}A)\big{(}I-PA_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA).$
Since
$A^{\frac{1}{2}}E_{\rm STG}=\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}=(I-F_{\rm TG})A^{\frac{1}{2}},$
we obtain
$\|E_{\rm STG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|(I-F_{\rm TG})\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm STG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}(I-F_{\rm TG})^{2}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
[…]
$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}(I_{r}-\Lambda_{r})^{2}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}.$
Due to
$Q_{1}\begin{pmatrix}I_{r}-\Lambda_{r}&0\\ 0&I_{n-r}\end{pmatrix}Q_{1}^{T}=I-F_{\rm TG}\succeq 0\quad\text{and}\quad\Lambda_{r}\succ 0,$
it follows that $\lambda(\Lambda_{r})\subset(0,1]$. Hence,
$\|E_{\rm STG}\|_{A}^{2}=\lambda_{\max}\big{(}(I_{r}-\Lambda_{r})^{2}\big{)}=\big{(}1-\lambda_{\min}(\Lambda_{r})\big{)}^{2}=\|E_{\rm TG}\|_{A}^{4},$
which yields
$\|E_{\rm STG}\|_{A}=\|E_{\rm TG}\|_{A}^{2}.$
This shows that there is no essential difference between the convergence analysis of Algorithm 1 and that of its symmetrized version.
Based on Lemma 3.1, we can derive the following convergence identity, which is the main result of this section.
Theorem 3.3.
Define
(3.5) $\widetilde{M}:=M+M^{T}-MAM^{T}$
and
(3.6) $\Pi_{A}:=PA_{\rm c}^{\dagger}P^{T}A.$
Under the assumptions of Lemma 3.1, it holds that
(3.7) $\|E_{\rm TG}\|_{A}=\sqrt{1-\sigma_{\rm TG}},$
where
(3.8) $\sigma_{\rm TG}=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
Proof.
Due to
$F_{\rm TG}=I-\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)},$
it follows that
$\lambda(F_{\rm TG})$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{)}$
$=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{)}$
$=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi\big{)}.$
According to the proof of Lemma 3.1, we deduce that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
Since $\Pi$ is an $L^{2}$-orthogonal projector and $\operatorname*{rank}(\Pi)=s$, there exists an orthogonal matrix $Q_{2}\in\mathbb{R}^{n\times n}$ such that
(3.9) $Q_{2}^{T}\Pi Q_{2}=\begin{pmatrix}I_{s}&0\\ 0&0\end{pmatrix}.$
Let
(3.10) $Q_{2}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}Q_{2}=\begin{pmatrix}X_{1}&X_{2}\\ X_{2}^{T}&X_{3}\end{pmatrix},$
where $X_{1}\in\mathbb{R}^{s\times s}$, $X_{2}\in\mathbb{R}^{s\times(n-s)}$, and $X_{3}\in\mathbb{R}^{(n-s)\times(n-s)}$. Then
$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
from which we deduce that $X_{3}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
Direct computation yields
$\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=Q_{2}\begin{pmatrix}I_{s}-X_{1}&-X_{2}\\ -X_{2}^{T}&I_{n-s}-X_{3}\end{pmatrix}Q_{2}^{T}.$
The positive semidefiniteness of $I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}$ implies that $\lambda_{\max}(X_{3})\leq 1$. Thus,
$\lambda_{n-r+1}(F_{\rm TG})=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi\big{)}=\lambda_{n-r+1}(X_{3}).$
Since
$(I-\Pi)A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)=Q_{2}\begin{pmatrix}0&0\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
we obtain
$\lambda_{n-r+1}(X_{3})$ $=\lambda_{n-r+s+1}\big{(}(I-\Pi)A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
$=\lambda_{n-r+s+1}\big{(}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)A^{\frac{1}{2}}\big{)}$
$=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
The identity (3.7) then follows by using (3.1). ∎
The identity (3.7) provides a direct way to analyze the convergence properties of Algorithm 1. A two-sided estimate for $\|E_{\rm TG}\|_{A}$ can be obtained by applying (3.7) and the following eigenvalue inequality; see, e.g., [12, Corollary 4.3.37].
Lemma 3.4.
Let $H\in\mathbb{C}^{n\times n}$ be Hermitian, and let $\{\mathbf{q}_{k}\}_{k=1}^{m}\subset\mathbb{C}^{n}\,(1\leq m\leq n)$ be a set of orthonormal vectors. Then, for any $i=1,\ldots,m$, it holds that
$\lambda_{i}(H)\leq\lambda_{i}(\check{H})\leq\lambda_{i+n-m}(H),$
where $\check{H}=\big{(}\mathbf{q}_{i}^{\ast}H\mathbf{q}_{j}\big{)}\in\mathbb{C}^{m\times m}$ and $(\cdot)^{\ast}$ denotes the conjugate transpose of a vector.
The following corollary presents a two-sided estimate for $\|E_{\rm TG}\|_{A}$.
Corollary 3.5.
Under the assumptions of Theorem 3.3, it holds that
(3.11) $\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}\leq\|E_{\rm TG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
Proof.
Since $I-\Pi$ is an $L^{2}$-orthogonal projector of rank $n-s$, there exists a matrix with orthonormal columns, $V_{n-s}\in\mathbb{R}^{n\times(n-s)}$, such that
$I-\Pi=V_{n-s}V_{n-s}^{T}.$
According to the proof of Theorem 3.3, it follows that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r+s$. Then
$\sigma_{\rm TG}$ $=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}\big{)}$
$=\lambda_{n-r+1}\big{(}V_{n-s}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}\big{)}.$
Using Lemma 3.4, we obtain
$\lambda_{n-r+1}(\widetilde{M}A)\leq\sigma_{\rm TG}\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
which, together with (3.7), yields the estimate (3.11). ∎
Remark 3.6.
From the relation
$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)},$
we deduce that
$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset(-\infty,1].$
In addition,
$\lambda(\widetilde{M}A)$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\subset[0,+\infty),$
where $\overline{M}$ is defined by (2.5). Hence, $\lambda(\widetilde{M}A)\subset[0,1]$.
4. An inexact variant of Algorithm 1 and its analysis
In this section, we analyze the convergence of an inexact variant of Algorithm 1, which is described by Algorithm 2. The output of its correction step, $\hat{\mathbf{e}}_{\rm c}$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$.
Algorithm 2  Inexact two-grid method.
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$
2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big{(}\mathbf{f}-A\mathbf{u}^{(1)}\big{)}$
3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$triangleright $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}:\mathbb{R}^{n_{\rm c}}\rightarrow\mathbb{R}^{n_{\rm c}}$
4:Prolongation: $\mathbf{u}_{\rm ITG}\leftarrow\mathbf{u}^{(1)}+P\hat{\mathbf{e}}_{\rm c}$
This section is divided into two subsections. The first one focuses on the convergence analysis of Algorithm 2 with linear coarse solvers; the second one is devoted to the case of general coarse solvers.
4.1. The linear case
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD approximation to $A_{\rm c}$. In this subsection, we study the convergence of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$.
We first prove an elementary lemma about the spectral relation between $A_{\rm c}$ and $B_{\rm c}$.
Lemma 4.1.
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be SPSD. Then, there exist positive numbers $c_{1}$ and $c_{2}$ such that
(4.1) $c_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}\leq c_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$
if and only if
(4.2) $\mathcal{R}(B_{\rm c})=\mathcal{R}(A_{\rm c}).$
Moreover, (4.1) is equivalent to the following inequality:
(4.3) $c_{2}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq c_{1}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}.$
Proof.
If (4.1) holds, then
$\mathcal{N}(B_{\rm c})=\mathcal{N}(A_{\rm c}),$
which is equivalent to (4.2). Conversely, if (4.2) holds, then
$\operatorname*{rank}(B_{\rm c})=\operatorname*{rank}(A_{\rm c})=s.$
Let
$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{2}\Sigma_{2}U_{2}^{T},$
where, for each $i\in\{1,2\}$, $U_{i}\in\mathbb{R}^{n_{\rm c}\times s}$ has orthonormal columns and $\Sigma_{i}\succ 0$ is an $s\times s$ diagonal matrix. The relation (4.2) implies that
$\mathcal{R}(U_{2})=\mathcal{R}(U_{1}).$
Then, there exists an orthogonal matrix $\Omega_{s}\in\mathbb{R}^{s\times s}$ such that $U_{2}=U_{1}\Omega_{s}$. Hence,
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Sigma_{1}U_{1}^{T}\mathbf{v}_{\rm c}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
$=\lambda_{\min}\big{(}\Sigma_{1}^{-1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}$
$=\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
Also, it holds that
$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
Therefore, the inequality (4.1) holds if $c_{1}\leq\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$ and $c_{2}\geq\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$.
We next prove the equivalence between (4.1) and (4.3). In view of (4.1), we have
$U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-c_{1}\Sigma_{1}\big{)}U_{1}^{T}=B_{\rm c}-c_{1}A_{\rm c}\succeq 0,$
from which we deduce that $\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succeq c_{1}\Sigma_{1}$ and thus $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$. Since
$c_{1}^{-1}A_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}=U_{1}\big{(}c_{1}^{-1}\Sigma_{1}^{-1}-\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{)}U_{1}^{T},$
it follows that $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$. Conversely, if $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$, then $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$, which leads to $B_{\rm c}\succeq c_{1}A_{\rm c}$.
Similarly, one can show that $c_{2}A_{\rm c}\succeq B_{\rm c}$ if and only if $B_{\rm c}^{\dagger}\succeq c_{2}^{-1}A_{\rm c}^{\dagger}$. The proof is completed. ∎
Based on Lemma 4.1 and its proof, we can further derive the following spectral relation.
Lemma 4.2.
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD matrix satisfying (4.2), and let
(4.4) $\alpha_{1}=\lambda_{n_{\rm c}-s+1}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}\quad\text{and}\quad\alpha_{2}=\lambda_{\max}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}.$
Define
$\widetilde{B}_{\rm c}^{\dagger}:=2B_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}A_{\rm c}B_{\rm c}^{\dagger}.$
If $\alpha_{2}<2$, then
(4.5) $\beta_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\beta_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}},$
where
(4.6a) $\beta_{1}$ =$=\begin{cases}(2-\alpha_{1})\alpha_{1}&\text{if $\alpha_{2}\leq 1$},\\ \min\big{\{}(2-\alpha_{1})\alpha_{1},\,(2-\alpha_{2})\alpha_{2}\big{\}}&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{2})\alpha_{2}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$},\end{cases}$
(4.6b) $\beta_{2}$ =$=\begin{cases}(2-\alpha_{2})\alpha_{2}&\text{if $\alpha_{2}\leq 1$},\\ 1&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{1})\alpha_{1}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$}.\end{cases}$
Proof.
Recall that
$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}U_{1}^{T}.$
Then
$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}$
and hence
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(\Omega_{s}^{T}U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
$=\lambda_{\min}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
Similarly, it holds that
$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
Note that
$\lambda_{\min}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{1}\quad\text{and}\quad\lambda_{\max}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{2},$
where $\alpha_{1}$ and $\alpha_{2}$ are given by (4.4). We then have
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{1}\quad\text{and}\quad\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{2},$
where $\beta_{1}$ and $\beta_{2}$ are given by (4.6a) and (4.6b), respectively. In addition,
$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{(}2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-\Sigma_{1}\big{)}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}U_{1}^{T},$
which, together with the condition $\alpha_{2}<2$ (or, equivalently, $2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succ\Sigma_{1}$), leads to
$\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}U_{1}^{T}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
Thus, the inequality (4.5) holds for all $\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$. ∎
Remark 4.3.
In the extreme case $B_{\rm c}=A_{\rm c}$, one has $\alpha_{1}=\alpha_{2}=1$ and $\beta_{1}=\beta_{2}=1$. We remark that $\alpha_{2}<2$ is a mild condition, which can be easily satisfied (e.g., one may simply scale $B_{\rm c}$ if necessary).
It is easy to see that the iteration matrix of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$ is
$E_{\rm ITG}=\big{(}I-PB_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA),$
which satisfies
$\mathbf{u}-\mathbf{u}_{\rm ITG}=E_{\rm ITG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}.$
Due to
$A^{\frac{1}{2}}E_{\rm ITG}=\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from Lemma 2.1 that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\|E_{\rm ITG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}$
with
$\|E_{\rm ITG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm ITG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}^{2}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(4.7) $F_{\rm ITG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Since $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$ and $\widetilde{B}_{\rm c}^{\dagger}\succeq 0$, it follows that $F_{\rm ITG}\succeq 0$. For any $\mathbf{v}\in\mathcal{N}(F_{\rm ITG})$, we have
$0=\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}=\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}+\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v},$
from which we deduce that
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\in\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
In view of (2.9), we have $\mathbf{v}\in\mathcal{N}(F_{\rm TG})$, which yields
$\mathcal{N}(F_{\rm ITG})\subseteq\mathcal{N}(F_{\rm TG}).$
Similarly, one can prove that $\mathcal{N}(F_{\rm TG})\subseteq\mathcal{N}(F_{\rm ITG})$. Thus,
(4.8) $\mathcal{N}(F_{\rm ITG})=\mathcal{N}(F_{\rm TG}),$
which implies that $\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$.
Note that
$\|E_{\rm ITG}\|_{A}<1$ $\Longleftrightarrow\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
$\Longleftrightarrow\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
$\Longleftrightarrow\|E_{\rm TG}\|_{A}<1.$
Hence, (2.10) is a sufficient and necessary condition for $\|E_{\rm ITG}\|_{A}<1$. In addition, from (3.4) and (4.8), we have
$\mathcal{N}(F_{\rm ITG})=\mathcal{N}(A).$
Similarly to the proof of Lemma 3.1, one can show that
(4.9) $\|E_{\rm ITG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm ITG})}.$
Let
$F_{i}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{i}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}$
for $i=1,2$. Using (4.5), (4.7), and (4.9), we obtain
(4.10) $\sqrt{1-\lambda_{n-r+1}(F_{2})}\leq\|E_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F_{1})}.$
To estimate the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$, we need the following eigenvalue inequalities; see, e.g., [12, Theorem 4.3.1].
Lemma 4.4.
Let $H_{1}\in\mathbb{C}^{n\times n}$ and $H_{2}\in\mathbb{C}^{n\times n}$ be Hermitian. For each $k=1,\ldots,n$, it holds that
(4.11a) $\lambda_{k}(H_{1}+H_{2})$ $\geq\lambda_{k-j+1}(H_{1})+\lambda_{j}(H_{2})\quad\forall\,j=1,\ldots,k,$
(4.11b) $\lambda_{k}(H_{1}+H_{2})$ $\leq\lambda_{k+\ell}(H_{1})+\lambda_{n-\ell}(H_{2})\quad\forall\,\ell=0,\ldots,n-k.$
In addition to (4.11a) and (4.11b), we need the following technical lemma.
Lemma 4.5.
Let $\Pi$, $\widetilde{M}$, and $\Pi_{A}$ be defined by (2.7), (3.5), and (3.6), respectively. Then
(4.12a) $\lambda_{\min}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=0,$
(4.12b) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=1-\delta_{\rm TG},$
where
(4.13) $\delta_{\rm TG}=\begin{cases}\lambda_{n-s+1}(\widetilde{M}A\Pi_{A})&\text{if $\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}=\{0\}$},\\[2.0pt] 0&\text{otherwise}.\end{cases}$
Proof.
The identity (4.12a) is obvious, because $\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\succeq 0$ is singular. Next, we focus on the proof of (4.12b).
Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]$, it follows from (3.10) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$. By (3.9) and (3.10), we have
$\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\Pi=Q_{2}\begin{pmatrix}I_{s}-X_{1}&0\\ -X_{2}^{T}&0\end{pmatrix}Q_{2}^{T}.$
Then
(4.14) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\Pi\big{)}$
$=1-\lambda_{\min}(X_{1}).$
Direct computation yields
$\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi=Q_{2}\begin{pmatrix}X_{1}&0\\ 0&0\end{pmatrix}Q_{2}^{T},$
where we have used the relations (3.9) and (3.10). It follows that
(4.15) $\operatorname*{rank}(X_{1})=\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}.$
We then prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
(4.16) $\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}=\{0\}.$
  • Recall that
    $\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\quad\text{and}\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
    If (4.16) holds, then
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}\subseteq\mathcal{N}(\Pi),$
    which, combined with the fact $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$, leads to
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
    Hence,
    $\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\operatorname*{rank}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\operatorname*{rank}(A_{\rm c})=s,$
    which, together with (4.15), leads to $\operatorname*{rank}(X_{1})=s$.
  • Conversely, if $\operatorname*{rank}(X_{1})=s$, then (4.16) holds. Otherwise, there must exist a vector $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(\Pi)$ such that
    $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\mathbf{v}=0,$
    i.e., $\mathbf{v}\in\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}\backslash\mathcal{N}(\Pi)$. This contradicts with the fact
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\mathcal{N}(\Pi),$
    which follows from the relations $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$ and
    $\operatorname*{rank}(\Pi)=s=\operatorname*{rank}(X_{1})=\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}.$
The above analysis shows that
[…]
Note that
(4.18) $\lambda\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\lambda\big{(}\widetilde{M}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{)}=\lambda(\widetilde{M}A\Pi_{A}).$
Combining (4.14), (4.17), and (4.18), we can arrive at the identity (4.12b). ∎
Remark 4.6.
The relation (4.16) will hold if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, since
$\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A)\quad\text{and}\quad\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\subseteq\mathcal{R}(A).$
Therefore, a sufficient condition for (4.16) to hold is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.6.
We are now ready to present a two-sided estimate for $\|E_{\rm ITG}\|_{A}$.
Theorem 4.7.
Under the assumptions of Algorithm 2 and Lemma 4.2, the inexact two-grid convergence factor $\|E_{\rm ITG}\|_{A}$ satisfies
(4.19) $\mathscr{L}(\beta_{2})\leq\|E_{\rm ITG}\|_{A}\leq\mathscr{U}(\beta_{1})$
with
$\mathscr{L}(\beta_{2})=\sqrt{1-\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}}$
and
(4.20) $\mathscr{U}(\beta_{1})=\sqrt{1-\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}},$
where $\widetilde{M}$, $\sigma_{\rm TG}$, $\beta_{1}$, $\beta_{2}$, and $\delta_{\rm TG}$ are given by (3.5), (3.8), (4.6a), (4.6b), and (4.13), respectively.
Proof.
In view of (4.10), one can get a two-sided estimate for $\|E_{\rm ITG}\|_{A}$ by bounding the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$.
A lower bound for $\lambda_{n-r+1}(F_{1})$: By (4.11a), we have
$\lambda_{n-r+1}(F_{1})$ $\geq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A),$
where we have used the facts (4.12a) and $\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}$. In addition, we have
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}(1-\beta_{1})A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{1}F_{\rm TG}\big{)}$
$\geq(1-\beta_{1})\lambda_{1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{n-r+1}(F_{\rm TG})$
$=\beta_{1}\sigma_{\rm TG}$
and
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{1})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{1})\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG}),$
where, in the last equality, we have used the fact (4.12b). Hence,
(4.21) $\lambda_{n-r+1}(F_{1})\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}.$
An upper bound for $\lambda_{n-r+1}(F_{2})$: Using (4.11b) and (4.12b), we obtain
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{2}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{2}\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}).$
Moreover,
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{2})\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG},$
where, in the last equality, we have used the fact (4.12a). Thus,
(4.22) $\lambda_{n-r+1}(F_{2})\leq\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}.$
The estimate (4.19) then follows by combining (4.10), (4.21), and (4.22). ∎
Remark 4.8.
By (4.11a) and (4.12b), we have
$1-\delta_{\rm TG}$ $=\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n}\big{(}F_{\rm TG}-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$=\sigma_{\rm TG}-\lambda_{n-r+1}(\widetilde{M}A),$
which gives
$\sigma_{\rm TG}\leq 1-\delta_{\rm TG}+\lambda_{n-r+1}(\widetilde{M}A).$
According to (4.6a) and (4.6b), we deduce that $0<\beta_{1}\leq\beta_{2}\leq 1$.
  • (i)
    If $\beta_{2}\rightarrow 0^{+}$, then $F_{i}\rightarrow A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}$ for $i=1,2$, which, together with (4.10), yields
    $\|E_{\rm ITG}\|_{A}\rightarrow\sqrt{1-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}}=\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
    This asymptotic behavior can also be observed from (4.19) as $\beta_{2}\rightarrow 0^{+}$.
  • (ii)
    In the special case $B_{\rm c}=A_{\rm c}$ (i.e., $\beta_{1}=\beta_{2}=1$), we have
    $\mathscr{L}(1)=\mathscr{U}(1)=\sqrt{1-\sigma_{\rm TG}},$
    which leads to the convergence identity (3.7).
Remark 4.9.
Since
$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]\quad\text{and}\quad\operatorname*{rank}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\leq\operatorname*{rank}\big{(}A^{\frac{1}{2}}\big{)}=r,$
it follows that
$\lambda_{n-r+1}(\widetilde{M}A)>0\Longleftrightarrow\operatorname*{rank}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=r\Longleftrightarrow\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
which will be satisfied if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$.
4.2. The general case
In this subsection, we present a convergence estimate for Algorithm 2 with general coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$, as described in the following theorem.
Theorem 4.10.
Under the assumptions of Algorithm 2, if
(4.23) $\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}$
for some $\varepsilon\in[0,1)$, then
(4.24) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\mathscr{U}(1-\varepsilon^{2})\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
where $\mathscr{U}(\cdot)$ is given by (4.20).
Proof.
The condition (4.23) reads
$2\big{(}A_{\rm c}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}\big{)}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
Note that $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
$\mathbf{r}_{\rm c}=P^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}\in\mathcal{R}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\mathcal{R}(A_{\rm c}).$
We then have
(4.25) $2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
From Algorithm 2, we have
$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}AP\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$\leq\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\big{(}I-(1-\varepsilon^{2})\Pi\big{)}A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)},$
where we have used the inequality (4.25). Due to
$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
and
$\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}},$
it follows that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}^{T}A^{\frac{1}{2}}(I-F)A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
with
$F=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+(1-\varepsilon^{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\bigg{(}1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}\bigg{)}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}^{2}.$
Similarly to the proof of Lemma 3.1, one can prove that
$\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}=\lambda_{n-r+1}(F).$
Thus,
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F)}\,\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}.$
The estimate (4.24) can be obtained by bounding $\lambda_{n-r+1}(F)$ from below. Indeed, it holds that
$\lambda_{n-r+1}(F)\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,(1-\varepsilon^{2})\sigma_{\rm TG},\,\sigma_{\rm TG}-\varepsilon^{2}(1-\delta_{\rm TG})\big{\}};$
see the proof of (4.21). This completes the proof. ∎
Remark 4.11.
From Algorithms 1 and 2, we have
$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\big{\|}P\big{(}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}\big{\|}_{A}=\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}},$
i.e., the perturbation bound of two-grid solution is equal to that of coarse solution. Therefore, the accuracy of two-grid solver can be controlled by that of coarse solver.
5. Conclusions
Two-grid convergence theory plays an important role in the design and analysis of multigrid methods. In this paper, we have established a new convergence theory for two-grid methods applied to SPSD linear systems. In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for the convergence factor of two-grid methods. More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers (both linear and general types are considered). A distinctive feature in comparison with the existing ones is that our theory does not require any additional assumptions on the system matrix, especially on its null space.
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (Grant No. 12401479), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241257), and the Scientific Research Foundation of the Southeast University (Grant No. 4007012324).

### v2

@@
-Two-grid convergence theory for symmetric positive semidefinite linear systems
+Convergence analysis of two-grid methods for symmetric positive semidefinite systems
 Abstract.
-This paper is devoted to the convergence theory of two-grid methods for symmetric positive semidefinite linear systems, with particular focus on the singular case.
-In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for characterizing the convergence factor of two-grid methods.
-More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers, including both linear and general cases.
-A key feature of our analysis is that it does not require any additional assumptions on the system matrix, especially on its null space.
+Two-grid theory plays a fundamental role in the design and analysis of multigrid methods.
+This paper is devoted to a new convergence analysis of two-grid methods for singular and symmetric positive semidefinite systems.
+Specifically, we derive a concise identity for characterizing the convergence factor of two-grid methods, with the Moore–Penrose inverse of coarse-grid matrix being used as a coarse solver.
+Furthermore, we present a convergence estimate for two-grid methods with approximate coarse solvers.
+Our new theory does not require any additional assumptions on the coefficient matrix, especially on its null space.
 Key words and phrases:
-two-grid methods, symmetric positive semidefinite matrices, singular systems, convergence analysis
+multigrid methods, two-grid methods, symmetric positive semidefinite matrices, null space, Moore–Penrose inverse
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-As a powerful solver with linear or near-linear computational complexity, multigrid has been widely used for the numerical solution of partial differential equations; see, e.g., [11, 4, 26, 27, 31].
+Multigrid is a powerful solver with linear or near-linear computational complexity that has been widely used for the numerical solution of partial differential equations; see, e.g., [12, 5, 30, 31, 36].
 The foundation of multigrid is a two-grid scheme, which consists of two complementary error-reduction processes: smoothing (or relaxation) and coarse-grid correction.
-In the correction step, a coarse-grid system with fewer unknowns needs to be solved.
+In general, the smoothing process is chosen as a simple iterative method (e.g., the weighted Jacobi and Gauss–Seidel methods), while these classical methods are commonly only effective at reducing high-frequency (or oscillatory) error.
+The remaining low-frequency (or smooth) error will be treated by the correction process, in which a coarse-grid system with fewer unknowns needs to be solved.
 In practice, it is often too costly to solve the coarse-grid system exactly.
-A typical strategy is to apply the two-grid scheme recursively in the correction steps, which will yield an algorithm with multilevel structure (i.e., multigrid method).
-For symmetric positive definite (SPD) linear systems, the convergence properties of two-grid and multigrid methods have been well studied; see, e.g., [29, 35, 30, 7, 8, 17, 31, 32, 33, 34].
-Many systems arising in mathematical models may only be semidefinite, such as discretized partial differential equations with Neumann boundary conditions (see, e.g., [1]) and Laplacian matrices of graphs (see, e.g., [24, 16]).
-In such cases, the system matrix, denoted by $A$, is singular and symmetric positive semidefinite (SPSD).
-Some convergence properties of classical iterative methods for SPSD systems can be found, e.g., in [13, 5, 14, 6, 9, 28, 10].
-The convergence of Schwarz methods for SPSD systems was studied in [21].
-Our focus here is on the convergence of two-grid methods for SPSD systems.
-In practice, two-grid convergence theory can be used to guide or motivate the design of multigrid methods; see, e.g., [3, 19, 18].
-Moreover, two-grid convergence estimates can be applied to the analysis of multigrid methods; see, e.g., [33, 34].
-In the SPD case, an identity for the convergence factor of two-grid methods was established in [8]; see [30, 36] for an abstract version.
-Ideally, one can derive a two-grid convergence identity for the semidefinite case without imposing any restrictions on the null space of $A$, which is denoted by $\mathcal{N}(A)$.
-Under several essential assumptions on local null spaces, a sharp convergence estimate for the successive subspace correction methods [29, 30] applied to SPSD systems was presented in [15]; see [28] for a simplified analysis.
-More specifically, [15] offers an identity for the seminorm of error propagation operator, despite it involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement.
-In [2], a classical two-grid convergence estimate for SPD systems [20, 23, 25] was extended to the semidefinite case.
-A two-grid convergence theory for singular systems was developed in [22], in which a restrictive condition on the coarse space and $\mathcal{N}(A)$ is required.
-Even in the special SPSD case, the two-grid estimates in [22] involve an extended prolongation operator, whose construction is based on $\mathcal{N}(A)$.
-A noteworthy contribution is the two-grid convergence identity appearing in [31, Theorem 5.3], which was established under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
-The aforementioned two-grid convergence theory and the associated assumptions rely largely on $\mathcal{N}(A)$, while the properties of $\mathcal{N}(A)$ may not be easily available.
-Consequently, it might be tough to apply those results or check the required conditions.
-Motivated by this observation, we establish a new convergence theory for two-grid methods applied to SPSD systems, which lifts the restrictions on $\mathcal{N}(A)$.
-Our main contribution is twofold.
-  • Firstly, we derive a succinct identity for the convergence factor of two-grid methods, provided that the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver (it is called the exact coarse solver).
-  • Secondly, we present some convergence estimates for two-grid methods with inexact coarse solvers, including both linear and general cases.
+A typical strategy is to apply the two-grid scheme recursively in the correction steps.
+The resulting methods (namely, multigrid methods) possess a multilevel structure.
+For symmetric positive definite (SPD) linear systems, the convergence of multigrid methods has been well studied; see, e.g., [33, 41, 34, 8, 9, 17, 36, 38, 39, 40, 37].
+Many important models or methods in mathematics may lead to symmetric positive semidefinite (SPSD) linear systems, such as partial differential equations with pure Neumann boundary conditions (see, e.g., [1]), generalized finite element methods (see, e.g., [26, 27, 28, 35]), and Laplacian matrices of graphs (see, e.g., [21, 25, 16]).
+Convergence of additive and multiplicative Schwarz methods for SPSD systems was studied in [22].
+Convergence theory of general stationary iterative methods for such systems can be found, e.g., in [13, 6, 14, 7, 10, 32, 11].
+Our focus here is on two-grid methods for SPSD systems.
+Two-grid theory plays a fundamental role in the design or analysis of multigrid methods; see, e.g., [3, 19, 18, 39, 40, 37].
+In the SPD case, an elegant identity for characterizing the convergence factor of two-grid methods was established in [9]; see [34, 42] for an abstract version.
+The identity is a powerful tool for analyzing the properties of two-grid methods; see, e.g., [9, 4, 38].
+Similarly, we expect to establish a convergence identity for two-grid methods applied to SPSD systems, without requiring any additional conditions on the null space of the coefficient matrix $A$, denoted by $\mathcal{N}(A)$.
+Under several essential conditions on local null spaces, a sharp convergence estimate for successive subspace correction methods [33, 34] applied to SPSD systems was presented in [15]; see [32] for a simplified analysis.
+More specifically, [15] offers an identity for the seminorm of error propagation operator, which involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement.
+In [2], a classical two-grid convergence estimate for SPD systems (see [20, 24, 29]) was extended to the semidefinite case.
+A two-grid convergence theory for singular systems was developed in [23], in which a condition on the coarse space and $\mathcal{N}(A)$ is required.
+Moreover, an extended prolongation operator, constructed from $\mathcal{N}(A)$, is needed.
+A noteworthy contribution is the two-grid identity in [36, Theorem 5.3], which was proved under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
+In some cases, however, the full information of $\mathcal{N}(A)$ may not be readily available.
+For example, as indicated in [26, page 52], the eigenfunctions associated with the zero eigenvalues are generally unknown in generalized finite element methods.
+In addition, most of the existing results focus on the case where the coarse-grid system is solved exactly, in the sense that a generalized inverse of the coarse-grid matrix is used as a coarse solver.
+For singular systems, more general coarse solvers are seldom considered in the literature.
+The difficulty may arise from the fact that the error propagation operator of coarse-grid correction fails to be a projector when the coarse-grid system is solved approximately.
+In this paper, we establish a new two-grid convergence theory for SPSD systems. Our main contribution is twofold:
+  • We derive a succinct identity for the convergence factor of two-grid methods, with the Moore–Penrose inverse of coarse-grid matrix being used as a coarse solver (for convenience, it is called an exact coarse solver);
+  • We present a convergence estimate for two-grid methods with inexact coarse solvers (various approximation strategies can be utilized to solve the coarse-grid system).
 It is worth mentioning that our theory has several distinctive features in comparison with the existing ones, as described below.
-  • It does not require any additional assumptions on $A$, especially on $\mathcal{N}(A)$.
-  • It yields an identity for the convergence factor of exact two-grid methods instead of an upper bound.
-  • It offers some convergence bounds for two-grid methods with inexact coarse solvers.
+  • It does not require any additional conditions on the coefficient matrix, especially on its null space.
+  • It gives an identity for the convergence factor of exact two-grid methods instead of an upper bound.
+  • It provides a convergence estimate for two-grid methods with inexact coarse solvers, that is, our analysis is not confined to the exact case.
 The rest of this paper is organized as follows.
@@
 In Section 3, we present a concise identity for the convergence factor of exact two-grid methods.
-In Section 4, we study the convergence of two-grid methods with approximate coarse solvers.
+In Section 4, we study the convergence of two-grid methods with approximate coarse solvers, and provide a numerical example to illustrate our theoretical estimate.
 In Section 5, we make some concluding remarks.
@@
 We start with some notation that will be used in the subsequent discussions.
-  • For two symmetric matrices $S_{1}$ and $S_{2}$, the symbol $S_{1}\succ S_{2}$ (resp., $S_{1}\succeq S_{2}$) means that $S_{1}-S_{2}$ is positive definite (resp., positive semidefinite).
+  • For a symmetric matrix $S$, $S\succ 0$ and $S\succeq 0$ mean that $S$ is positive definite and positive semidefinite, respectively.
   • $I_{n}$ denotes the $n\times n$ identity matrix (or $I$ when its size is clear from context).
-  • $\mathcal{R}(\cdot)$ denotes the range (or column space) of a matrix.
-  • $\mathcal{N}(\cdot)$ denotes the null space of a matrix.
-  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix, respectively.
-  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix.
+  • $\mathcal{R}(\cdot)$ denotes the column space (or range) of a matrix.
+  • $\mathcal{N}(\cdot)$ denotes the null space (or kernel) of a matrix.
+  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix with real eigenvalues, respectively.
+  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix with real eigenvalues.
   • $\lambda(\cdot)$ denotes the spectrum of a matrix.
@@
 That is, (2.1) defines a seminorm on $\mathbb{R}^{n\times n}$.
-The following lemma gives two simple but useful properties of the $A$-seminorm.
+The following lemma gives two elementary but useful properties of the $A$-seminorm.
 Lemma 2.1.
@@
 and
-(2.3) $\|X\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
+[…]
 Proof.
@@
 $\mathcal{N}(A)=\mathcal{N}\big(A^{\frac{1}{2}}\big)\quad\text{and}\quad\mathcal{R}(A)=\mathcal{R}\big(A^{\frac{1}{2}}\big).$
-It can be readily seen from (2.1) that (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
+It can be readily seen from (2.1) that the inequality (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
 $\|X\mathbf{v}\|_{A}=\big\|A^{\frac{1}{2}}X\mathbf{v}\big\|_{2}=\big\|YA^{\frac{1}{2}}\mathbf{v}\big\|_{2}=0,$
@@
 $\|X\|_{A}$ $=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big\|A^{\frac{1}{2}}X\mathbf{v}\big\|_{2}}{\big\|A^{\frac{1}{2}}\mathbf{v}\big\|_{2}}=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big\|YA^{\frac{1}{2}}\mathbf{v}\big\|_{2}}{\big\|A^{\frac{1}{2}}\mathbf{v}\big\|_{2}}$
-$=\sup_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}},$
+[…]
 which yields the expression (2.3). ∎
-Consider the solution of linear systems of the form
+Consider solving the linear system
 (2.4) $A\mathbf{u}=\mathbf{f},$
-where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$, $A\succeq 0$, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$.
-When $A$ is singular (namely, $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$.
-We aim to solve (2.4) using a two-grid method, the output of which is an approximation to a solution to (2.4).
-Two very basic assumptions involved in the two-grid analysis are made below.
+where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$ is SPSD, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$ (that is, $\mathbf{f}$ is orthogonal to $\mathcal{N}(A)$).
+When $A$ is singular (i.e., $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$.
+We will analyze the convergence of two-grid methods for solving (2.4).
+Two basic assumptions involved in the analysis of two-grid methods are made below.
+  • (A1):
     Let $M$ be an $n\times n$ matrix such that $\|I-MA\|_{A}\leq 1$.
+  • (A2):
     Let $P$ be an $n\times n_{\rm c}$ prolongation (or interpolation) matrix such that the Galerkin coarse-grid matrix $A_{\rm c}:=P^{T}AP$ is nonzero, where $n_{\rm c}\,(<n)$ is the number of coarse variables.
 Remark 2.2.
-We remark that $M$ is not necessarily nonsingular.
-Moreover, measured in $\|\cdot\|_{A}$, the smoothing iteration matrix in Algorithm 1 below (i.e., $I-MA$) is not required to be a strict contraction.
+In view of (A1) and (A2), $M$ is not necessarily nonsingular, and $P$ is not necessarily of full column rank.
+In addition, measured in $\|\cdot\|_{A}$, the smoothing iteration in Algorithm 1 below is not required to be a strict contraction.
 Due to
@@
 it follows from (2.3) that
-$\|I-MA\|_{A}^{2}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}^{2}}{\|\mathbf{v}\|_{2}^{2}}=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
+[…]
 where
 (2.5) $\overline{M}:=M+M^{T}-M^{T}AM.$
-Thus, $\|I-MA\|_{A}\leq 1$ if and only if $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$, or, equivalently, $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathbb{R}^{n}\backslash\{0\}$ (i.e., $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$), because $\mathbb{R}^{n}$ is the direct sum of $\mathcal{R}(A)$ and $\mathcal{N}(A)$, denoted by $\mathbb{R}^{n}=\mathcal{R}(A)\oplus\mathcal{N}(A)$.
+[…]
 With the above assumptions, a two-grid method for solving (2.4) is described by Algorithm 1, in which the Moore–Penrose inverse $A_{\rm c}^{\dagger}$ is used as a coarse solver.
@@
 Algorithm 1 Exact two-grid method.
-1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$triangleright $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
+1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big(\mathbf{f}-A\mathbf{u}^{(0)}\big)$ $\triangleright$ $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
 2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big(\mathbf{f}-A\mathbf{u}^{(1)}\big)$
@@
 Remark 2.3.
-The output of Algorithm 1 converges to a particular solution to (2.4) as $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}$ tends to zero.
-Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}=A\mathbf{u}=\mathbf{f}$.
+In Algorithm 1, the coarse-grid system to be solved reads
+(2.6) $A_{\rm c}\mathbf{x}_{\rm c}=\mathbf{r}_{\rm c}.$
+It is easy to see that $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$ is a solution to (2.6), because $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
+$\mathbf{r}_{\rm c}=P^{T}\big(\mathbf{f}-A\mathbf{u}^{(1)}\big)=P^{T}A\big(\mathbf{u}-\mathbf{u}^{(1)}\big)\in\mathcal{R}\big(P^{T}A^{\frac{1}{2}}\big)=\mathcal{R}(A_{\rm c}).$
 From Algorithm 1, we have
@@
 It can be easily verified that
-$\Pi^{T}=\Pi=\Pi^{2},\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)},\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}.$
-Thus, $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}$ onto $\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}$. Since
+$\Pi^{2}=\Pi=\Pi^{T},\quad\mathcal{N}(\Pi)=\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big),\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big(A^{\frac{1}{2}}P\big),$
+i.e., $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big)$ onto $\mathcal{R}\big(A^{\frac{1}{2}}P\big)$. Since
 $A^{\frac{1}{2}}E_{\rm TG}=(I-\Pi)\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)A^{\frac{1}{2}},$
 it follows from Lemma 2.1 and (2.7) that
-$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
+$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
 where $\|E_{\rm TG}\|_{A}$, called the convergence factor of Algorithm 1, is given by
@@
 Then
-$\|E_{\rm TG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
-$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
+[…]
+[…]
 where
 (2.10) $F_{\rm TG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big).$
-Hence, it holds that $\|E_{\rm TG}\|_{A}<1$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
+[…]
+Remark 2.4.
+If $\|E_{\rm TG}\|_{A}<1$ and Algorithm 1 is carried out $\nu$ times iteratively (the corresponding output is denoted by $\mathbf{u}_{\rm TG}^{(\nu)}$), then $\|\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\|_{A}\to 0$ as $\nu\to\infty$, in which case $\big\{\mathbf{u}_{\rm TG}^{(\nu)}\big\}_{\nu=1}^{\infty}$ converge to a particular solution to (2.4).
+Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}^{(\nu)}=A\mathbf{u}=\mathbf{f}$.
 The following lemma provides a sufficient and necessary condition (namely, (2.11)) for $\|E_{\rm TG}\|_{A}<1$.
 It is worth pointing out that (2.11) is not really a restriction on $\mathcal{N}(A)$.
-In view of (2.10), a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$, which in fact can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.6.
-Throughout this paper, we always assume that (2.10) holds.
+In fact, we can deduce from (2.11) that a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$.
+This condition can be satisfied by the conventional weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.7 below.
+Throughout this paper, we always assume that (A1), (A2), and (2.11) hold.
 Lemma 2.5.
-Under the assumptions of Algorithm 1, $\|E_{\rm TG}\|_{A}<1$ if and only if
-(2.10) $\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
+The convergence factor $\|E_{\rm TG}\|_{A}$ is strictly less than 11 if and only if
+(2.11) $\mathcal{N}\big(A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big)\cap\mathcal{N}\big(P^{T}(I-AM)A^{\frac{1}{2}}\big)=\mathcal{N}(A),$
+where $\overline{M}$ is defined by (2.5).
 Proof.
 Observe first that
-$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq 0}+\underbrace{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}_{\geq 0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
+$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq\,0}+\underbrace{\mathbf{v}^{T}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\mathbf{v}}_{\geq\,0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
 and
@@
 $\mathbf{x}_{1}^{T}F_{\rm TG}\mathbf{x}_{1}=\mathbf{x}_{1}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}_{1}+\mathbf{x}_{1}^{T}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\mathbf{x}_{1}=0,$
-which contradicts with the condition $\|E_{\rm TG}\|_{A}<1$, or, equivalently, $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
-(2.10)$\Rightarrow\|E_{\rm TG}\|_{A}<1$”: Conversely, if (2.10) holds, then $\|E_{\rm TG}\|_{A}<1$.
-Otherwise, there must exist a vector $\mathbf{y}\in\mathcal{R}(A)\backslash\{0\}$ such that $\mathbf{y}^{T}F_{\rm TG}\mathbf{y}=0$.
-Then
+[…]
+[…]
 $\mathbf{y}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{y}=0\quad\text{and}\quad\mathbf{y}^{T}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\mathbf{y}=0,$
@@
 Corollary 2.6.
-Under the assumptions of Algorithm 1, if $\overline{M}\succeq 0$ and
+Let $\overline{M}$ be defined by (2.5). If $\overline{M}\succeq 0$ and
 (2.12) $\mathcal{N}(\overline{M})\cap\mathcal{R}(A)=\{0\},$
@@
 Remark 2.7.
-From Corollary 2.5, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, provided that the diagonal entries of $A$ are all nonzero.
 Let
@@
 If some diagonal entry of $D$ is zero, the corresponding row and column in $A$ are both zero, in which case one may solve a reduced version of (2.4) obtained by deleting the zero row and column.
+From Corollary 2.6, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers.
   • If $M=\omega D^{-1}$ with $0<\omega<\frac{2}{\lambda_{\max}(D^{-1}A)}$, then
     $\overline{M}=\omega^{2}D^{-1}\big(2\omega^{-1}D-A\big)D^{-1}\succ 0.$
-Another typical choice is $M=(D+L)^{-1}$, in which case one has
+  • If $M=(D+L)^{-1}$, then
     $\overline{M}=(D+L)^{-T}D(D+L)^{-1}\succ 0.$
+In both cases, it holds that $\|E_{\rm TG}\|_{A}<1$.
 3. Convergence analysis of Algorithm 1
 In this section, we present a succinct identity for characterizing the convergence factor $\|E_{\rm TG}\|_{A}$.
-Define
-$r:=\operatorname*{rank}(A)\quad\text{and}\quad s:=\operatorname*{rank}(A_{\rm c}).$
+Let
+$r=\operatorname*{rank}(A)\quad\text{and}\quad s=\operatorname*{rank}(A_{\rm c}).$
 Since $A_{\rm c}=P^{T}AP$, it follows that $s\leq r$. In the extreme case $s=r$, we have
@@
 $(I-\Pi)\mathbf{v}=0\quad\forall\,\mathbf{v}\in\mathcal{R}(A),$
-which, together with (2.8), yields $\|E_{\rm TG}\|_{A}=0$. Such an extreme case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
+which, together with (2.9), yields $\|E_{\rm TG}\|_{A}=0$. However, such a case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
 We first prove a technical lemma, which gives a reformulation of the convergence factor $\|E_{\rm TG}\|_{A}$.
 Lemma 3.1.
-Under the assumptions of Algorithm 1 and the condition (2.10), one has
+The convergence factor $\|E_{\rm TG}\|_{A}$ can be reformulated as
 (3.1) $\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm TG})},$
-where $F_{\rm TG}$ is given by (2.9).
+where $r=\operatorname*{rank}(A)$ and $F_{\rm TG}$ is given by (2.10).
 Proof.
@@
 Combining (3.2) and (3.3), we obtain
-(3.4) $\mathcal{N}(F_{\rm TG})=\mathcal{N}(A)$
+$\mathcal{N}(F_{\rm TG})=\mathcal{N}(A)$
 and hence
@@
 Thus,
-$\|E_{\rm TG}\|_{A}^{2}$ $=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
 […]
-$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\Lambda_{r}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}$
+$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}F_{\rm TG}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}$
+[…]
 $=1-\lambda_{\min}(\Lambda_{r})$
@@
 Then
-$\|E_{\rm STG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}(I-F_{\rm TG})^{2}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
 […]
-$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}(I_{r}-\Lambda_{r})^{2}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}.$
+$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}(I-F_{\rm TG})^{2}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}$
+[…]
 Due to
@@
 This shows that there is no essential difference between the convergence analysis of Algorithm 1 and that of its symmetrized version.
-Based on Lemma 3.1, we can derive the following convergence identity, which is the main result of this section.
+Based on (3.1) and its proof, we can further derive the following identity, which is the main result of this section.
 Theorem 3.3.
@@
 (3.5) $\Pi_{A}:=PA_{\rm c}^{\dagger}P^{T}A.$
-Under the assumptions of Lemma 3.1, it holds that
-(3.7) $\|E_{\rm TG}\|_{A}=\sqrt{1-\sigma_{\rm TG}},$
-where
-(3.8) $\sigma_{\rm TG}=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
+Then, it holds that
+(3.6) $\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{n-r+s+1}\big(\widetilde{M}A(I-\Pi_{A})\big)},$
+where $r=\operatorname*{rank}(A)$ and $s=\operatorname*{rank}(A_{\rm c})$.
 Proof.
@@
 $=\lambda\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi\big).$
-According to the proof of Lemma 3.1, we deduce that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
 Since $\Pi$ is an $L^{2}$-orthogonal projector and $\operatorname*{rank}(\Pi)=s$, there exists an orthogonal matrix $Q_{2}\in\mathbb{R}^{n\times n}$ such that
@@
 where $X_{1}\in\mathbb{R}^{s\times s}$, $X_{2}\in\mathbb{R}^{s\times(n-s)}$, and $X_{3}\in\mathbb{R}^{(n-s)\times(n-s)}$. Then
-$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
-from which we deduce that $X_{3}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
+$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T}.$
+Note that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and nrn-r zero eigenvalues; see the proof of (3.1).
+We conclude that $X_{3}$ has rsr-s positive eigenvalues and nrn-r zero eigenvalues.
 Direct computation yields
@@
 The identity (3.6) then follows by using (3.1). ∎
-The identity (3.7) provides a direct way to analyze the convergence properties of Algorithm 1.
-A two-sided estimate for $\|E_{\rm TG}\|_{A}$ can be obtained by applying (3.7) and the following eigenvalue inequality; see, e.g., [12, Corollary 4.3.37].
-Lemma 3.4.
-Let $H\in\mathbb{C}^{n\times n}$ be Hermitian, and let $\{\mathbf{q}_{k}\}_{k=1}^{m}\subset\mathbb{C}^{n}\,(1\leq m\leq n)$ be a set of orthonormal vectors. Then, for any $i=1,\ldots,m$, it holds that
-$\lambda_{i}(H)\leq\lambda_{i}(\check{H})\leq\lambda_{i+n-m}(H),$
-where $\check{H}=\big{(}\mathbf{q}_{i}^{\ast}H\mathbf{q}_{j}\big{)}\in\mathbb{C}^{m\times m}$ and $(\cdot)^{\ast}$ denotes the conjugate transpose of a vector.
-The following corollary presents a two-sided estimate for $\|E_{\rm TG}\|_{A}$.
-Corollary 3.5.
-Under the assumptions of Theorem 3.3, it holds that
-(3.11) $\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}\leq\|E_{\rm TG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
-Proof.
-Since $I-\Pi$ is an $L^{2}$-orthogonal projector of rank $n-s$, there exists a matrix with orthonormal columns, $V_{n-s}\in\mathbb{R}^{n\times(n-s)}$, such that
-$I-\Pi=V_{n-s}V_{n-s}^{T}.$
-According to the proof of Theorem 3.3, it follows that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r+s$. Then
-$\sigma_{\rm TG}$ $=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}$
-$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
-$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}\big{)}$
-$=\lambda_{n-r+1}\big{(}V_{n-s}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}\big{)}.$
-Using Lemma 3.4, we obtain
-$\lambda_{n-r+1}(\widetilde{M}A)\leq\sigma_{\rm TG}\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
-which, together with (3.7), yields the estimate (3.11). ∎
 Remark 3.4.
-From the relation
-$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)},$
-we deduce that
-$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset(-\infty,1].$
+In view of (3.4), we have
+$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\succeq 0,$
+which leads to
+$\lambda_{\max}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)\leq 1.$
 In addition,
-$\lambda(\widetilde{M}A)$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{)}$
+$\lambda\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)$ $=\lambda\big(I-\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\big)$
 $=\lambda\big(I-\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big)$
 $=\lambda\big(A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big)\subset[0,+\infty),$
-where $\overline{M}$ is defined by (2.5). Hence, $\lambda(\widetilde{M}A)\subset[0,1]$.
+where $\overline{M}$ is defined by (2.5). Hence, $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}$ is an SPSD matrix with eigenvalues contained in $[0,1]$. Since
+$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}-\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)^{\frac{1}{2}}(I-\Pi)\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)^{\frac{1}{2}}\succeq 0,$
+it follows that
+$\lambda_{n-r+s+1}\big(\widetilde{M}A(I-\Pi_{A})\big)=\lambda_{n-r+s+1}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big)\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
+which, combined with (3.6), gives
+(3.9) $\|E_{\rm TG}\|_{A}\geq\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}.$
+In the special case where $A\succ 0$ and $\operatorname*{rank}(P)=n_{\rm c}$ (i.e., $r=n$ and $s=n_{\rm c}$), (3.9) reduces to
+(3.10) $\|E_{\rm TG}\|_{A}\geq\sqrt{1-\lambda_{n_{\rm c}+1}(\widetilde{M}A)}.$
+As proved in [4, Lemma 1], the lower bound in (3.10) is attainable.
+Furthermore, (3.10) and its sharpness still hold when $A\succeq 0$, $\mathcal{N}(A)\subset\mathcal{R}(P)$, and $\operatorname*{rank}(P)=n_{\rm c}$; see [36, Theorem 5.6].
+Using (3.6), we can prove the following corollary.
+Corollary 3.5.
+Let
+$d_{1}=\dim\big(\mathcal{R}(P)\cap\mathcal{N}(A)\big)\quad\text{and}\quad d_{2}=\dim\big(\mathcal{R}(P)+\mathcal{N}(A)\big).$
+Then, it holds that
+(3.11a) $\|E_{\rm TG}\|_{A}$ $=\sqrt{1-\lambda_{n-r+\operatorname*{rank}(P)-d_{1}+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}$
+(3.11b) $=\sqrt{1-\lambda_{d_{2}+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}.$
+Proof.
+Assume that $\{\mathbf{p}_{1},\ldots,\mathbf{p}_{d_{1}}\}$ is a basis for $\mathcal{R}(P)\cap\mathcal{N}(A)$.
+Such a set can be expanded as a basis for $\mathcal{R}(P)$, which is denoted by $\big\{\mathbf{p}_{1},\ldots,\mathbf{p}_{d_{1}},\mathbf{p}_{d_{1}+1},\ldots,\mathbf{p}_{\operatorname*{rank}(P)}\big\}$.
+Let
+$P=\big(\mathbf{p}_{1},\ldots,\mathbf{p}_{\operatorname*{rank}(P)}\big)W,$
+where $W\in\mathbb{R}^{\operatorname*{rank}(P)\times n_{\rm c}}$ is of full row rank. Then
+$AP=\big(0,\ldots,0,A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)W.$
+On the one hand,
+$\operatorname*{rank}(AP)\leq\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big).$
+On the other hand,
+$\operatorname*{rank}(AP)$ $\geq\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big)+\operatorname*{rank}(W)-\operatorname*{rank}(P)$
+$=\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big).$
+Note that the set $\big\{A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big\}$ is linearly independent. We then have
+$\operatorname*{rank}(AP)=\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big)=\operatorname*{rank}(P)-d_{1}.$
+Due to
+$\mathcal{N}(A_{\rm c})=\mathcal{N}(A^{\frac{1}{2}}P)=\mathcal{N}(AP),$
+it follows that
+$s=\operatorname*{rank}(A_{\rm c})=\operatorname*{rank}(AP)=\operatorname*{rank}(P)-d_{1},$
+which, together with (3.6), yields the identity (3.11a).
+The identity (3.11b) follows from (3.11a) and the fact
+$d_{1}=\operatorname*{rank}(P)+n-r-d_{2},$
+because
+$\dim\big(\mathcal{R}(P)\big)+\dim\big(\mathcal{N}(A)\big)=\dim\big(\mathcal{R}(P)+\mathcal{N}(A)\big)+\dim\big(\mathcal{R}(P)\cap\mathcal{N}(A)\big).$
+This completes the proof. ∎
+In particular, if $\mathcal{N}(A)\subset\mathcal{R}(P)$, then we have the following result, which provides an alternative to the convergence identity proved in [36, Theorem 5.3].
+Corollary 3.6.
+If $\mathcal{N}(A)\subset\mathcal{R}(P)$, then
+$\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{\operatorname*{rank}(P)+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}.$
+Proof.
+If $\mathcal{N}(A)\subset\mathcal{R}(P)$, then
+$d_{1}=n-r\quad\text{and}\quad d_{2}=\operatorname*{rank}(P).$
+The desired result then follows immediately from (3.11a) or (3.11b). ∎
 4. An inexact variant of Algorithm 1 and its analysis
-In this section, we analyze the convergence of an inexact variant of Algorithm 1, which is described by Algorithm 2.
-The output of its correction step, $\hat{\mathbf{e}}_{\rm c}$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$.
+In this section, we are devoted to analyzing the convergence of an inexact variant of Algorithm 1 (that is, the coarse-grid system (2.6) is solved approximately), which is described by Algorithm 2 below.
+The output of its correction step, $\hat{\mathbf{e}}_{\rm c}=\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$.
+Here, $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}$ is a general mapping from $\mathbb{R}^{n_{\rm c}}$ to $\mathbb{R}^{n_{\rm c}}$.
 Algorithm 2 Inexact two-grid method.
@@
 2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big(\mathbf{f}-A\mathbf{u}^{(1)}\big)$
-3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$triangleright $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}:\mathbb{R}^{n_{\rm c}}\rightarrow\mathbb{R}^{n_{\rm c}}$
+3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$
 4:Prolongation: $\mathbf{u}_{\rm ITG}\leftarrow\mathbf{u}^{(1)}+P\hat{\mathbf{e}}_{\rm c}$
-This section is divided into two subsections.
-The first one focuses on the convergence analysis of Algorithm 2 with linear coarse solvers; the second one is devoted to the case of general coarse solvers.
-4.1. The linear case
-Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD approximation to $A_{\rm c}$.
-In this subsection, we study the convergence of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$.
-We first prove an elementary lemma about the spectral relation between $A_{\rm c}$ and $B_{\rm c}$.
-Lemma 4.1.
-Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be SPSD. Then, there exist positive numbers $c_{1}$ and $c_{2}$ such that
-(4.1) $c_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}\leq c_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$
-if and only if
-(4.2) $\mathcal{R}(B_{\rm c})=\mathcal{R}(A_{\rm c}).$
-Moreover, (4.1) is equivalent to the following inequality:
-(4.3) $c_{2}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq c_{1}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}.$
-Proof.
-If (4.1) holds, then
-$\mathcal{N}(B_{\rm c})=\mathcal{N}(A_{\rm c}),$
-which is equivalent to (4.2). Conversely, if (4.2) holds, then
-$\operatorname*{rank}(B_{\rm c})=\operatorname*{rank}(A_{\rm c})=s.$
-Let
-$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{2}\Sigma_{2}U_{2}^{T},$
-where, for each $i\in\{1,2\}$, $U_{i}\in\mathbb{R}^{n_{\rm c}\times s}$ has orthonormal columns and $\Sigma_{i}\succ 0$ is an $s\times s$ diagonal matrix. The relation (4.2) implies that
-$\mathcal{R}(U_{2})=\mathcal{R}(U_{1}).$
-Then, there exists an orthogonal matrix $\Omega_{s}\in\mathbb{R}^{s\times s}$ such that $U_{2}=U_{1}\Omega_{s}$. Hence,
-$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Sigma_{1}U_{1}^{T}\mathbf{v}_{\rm c}}$
-$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
-$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
-$=\lambda_{\min}\big{(}\Sigma_{1}^{-1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}$
-$=\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
-Also, it holds that
-$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
-Therefore, the inequality (4.1) holds if $c_{1}\leq\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$ and $c_{2}\geq\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$.
-We next prove the equivalence between (4.1) and (4.3). In view of (4.1), we have
-$U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-c_{1}\Sigma_{1}\big{)}U_{1}^{T}=B_{\rm c}-c_{1}A_{\rm c}\succeq 0,$
-from which we deduce that $\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succeq c_{1}\Sigma_{1}$ and thus $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$. Since
-$c_{1}^{-1}A_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}=U_{1}\big{(}c_{1}^{-1}\Sigma_{1}^{-1}-\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{)}U_{1}^{T},$
-it follows that $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$.
-Conversely, if $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$, then $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$, which leads to $B_{\rm c}\succeq c_{1}A_{\rm c}$.
-Similarly, one can show that $c_{2}A_{\rm c}\succeq B_{\rm c}$ if and only if $B_{\rm c}^{\dagger}\succeq c_{2}^{-1}A_{\rm c}^{\dagger}$. The proof is completed. ∎
-Based on Lemma 4.1 and its proof, we can further derive the following spectral relation.
-Lemma 4.2.
-Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD matrix satisfying (4.2), and let
-(4.4) $\alpha_{1}=\lambda_{n_{\rm c}-s+1}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}\quad\text{and}\quad\alpha_{2}=\lambda_{\max}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}.$
-Define
-$\widetilde{B}_{\rm c}^{\dagger}:=2B_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}A_{\rm c}B_{\rm c}^{\dagger}.$
-If $\alpha_{2}<2$, then
-(4.5) $\beta_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\beta_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}},$
-where
-(4.6a) $\beta_{1}$ =$=\begin{cases}(2-\alpha_{1})\alpha_{1}&\text{if $\alpha_{2}\leq 1$},\\ \min\big{\{}(2-\alpha_{1})\alpha_{1},\,(2-\alpha_{2})\alpha_{2}\big{\}}&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{2})\alpha_{2}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$},\end{cases}$
-(4.6b) $\beta_{2}$ =$=\begin{cases}(2-\alpha_{2})\alpha_{2}&\text{if $\alpha_{2}\leq 1$},\\ 1&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{1})\alpha_{1}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$}.\end{cases}$
-Proof.
-Recall that
-$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}U_{1}^{T}.$
-Then
-$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}$
-and hence
-$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}$
-$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(\Omega_{s}^{T}U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
-$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
-$=\lambda_{\min}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
-Similarly, it holds that
-$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
-Note that
-$\lambda_{\min}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{1}\quad\text{and}\quad\lambda_{\max}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{2},$
-where $\alpha_{1}$ and $\alpha_{2}$ are given by (4.4). We then have
-$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{1}\quad\text{and}\quad\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{2},$
-where $\beta_{1}$ and $\beta_{2}$ are given by (4.6a) and (4.6b), respectively. In addition,
-$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{(}2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-\Sigma_{1}\big{)}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}U_{1}^{T},$
-which, together with the condition $\alpha_{2}<2$ (or, equivalently, $2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succ\Sigma_{1}$), leads to
-$\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}U_{1}^{T}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
-Thus, the inequality (4.5) holds for all $\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$. ∎
-Remark 4.3.
-In the extreme case $B_{\rm c}=A_{\rm c}$, one has $\alpha_{1}=\alpha_{2}=1$ and $\beta_{1}=\beta_{2}=1$.
-We remark that $\alpha_{2}<2$ is a mild condition, which can be easily satisfied (e.g., one may simply scale $B_{\rm c}$ if necessary).
-It is easy to see that the iteration matrix of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$ is
-$E_{\rm ITG}=\big{(}I-PB_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA),$
-which satisfies
-$\mathbf{u}-\mathbf{u}_{\rm ITG}=E_{\rm ITG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}.$
-Due to
-$A^{\frac{1}{2}}E_{\rm ITG}=\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
-it follows from Lemma 2.1 that
-$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\|E_{\rm ITG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}$
-with
-$\|E_{\rm ITG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}}{\|\mathbf{v}\|_{2}}.$
-Then
-$\|E_{\rm ITG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}^{2}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
-$=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
-$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
-where
-(4.7) $F_{\rm ITG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
-Since $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$ and $\widetilde{B}_{\rm c}^{\dagger}\succeq 0$, it follows that $F_{\rm ITG}\succeq 0$.
-For any $\mathbf{v}\in\mathcal{N}(F_{\rm ITG})$, we have
-$0=\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}=\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}+\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v},$
-from which we deduce that
-$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\in\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
-In view of (2.9), we have $\mathbf{v}\in\mathcal{N}(F_{\rm TG})$, which yields
-$\mathcal{N}(F_{\rm ITG})\subseteq\mathcal{N}(F_{\rm TG}).$
-Similarly, one can prove that $\mathcal{N}(F_{\rm TG})\subseteq\mathcal{N}(F_{\rm ITG})$. Thus,
-(4.8) $\mathcal{N}(F_{\rm ITG})=\mathcal{N}(F_{\rm TG}),$
-which implies that $\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$.
-Note that
-$\|E_{\rm ITG}\|_{A}<1$ $\Longleftrightarrow\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
-$\Longleftrightarrow\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
-$\Longleftrightarrow\|E_{\rm TG}\|_{A}<1.$
-Hence, (2.10) is a sufficient and necessary condition for $\|E_{\rm ITG}\|_{A}<1$. In addition, from (3.4) and (4.8), we have
-$\mathcal{N}(F_{\rm ITG})=\mathcal{N}(A).$
-Similarly to the proof of Lemma 3.1, one can show that
-(4.9) $\|E_{\rm ITG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm ITG})}.$
-Let
-$F_{i}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{i}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}$
-for $i=1,2$. Using (4.5), (4.7), and (4.9), we obtain
-(4.10) $\sqrt{1-\lambda_{n-r+1}(F_{2})}\leq\|E_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F_{1})}.$
-To estimate the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$, we need the following eigenvalue inequalities; see, e.g., [12, Theorem 4.3.1].
-Lemma 4.4.
-Let $H_{1}\in\mathbb{C}^{n\times n}$ and $H_{2}\in\mathbb{C}^{n\times n}$ be Hermitian. For each $k=1,\ldots,n$, it holds that
-(4.11a) $\lambda_{k}(H_{1}+H_{2})$ $\geq\lambda_{k-j+1}(H_{1})+\lambda_{j}(H_{2})\quad\forall\,j=1,\ldots,k,$
-(4.11b) $\lambda_{k}(H_{1}+H_{2})$ $\leq\lambda_{k+\ell}(H_{1})+\lambda_{n-\ell}(H_{2})\quad\forall\,\ell=0,\ldots,n-k.$
-In addition to (4.11a) and (4.11b), we need the following technical lemma.
+This section will be divided into two parts: the first part focuses on a theoretical estimate of $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}$; the second part provides a numerical example to illustrate our estimate.
+4.1. Convergence analysis
+To analyze the convergence of Algorithm 2, we need the following technical lemma.
 Lemma 4.1.
 Let $\Pi$, $\widetilde{M}$, and $\Pi_{A}$ be defined by (2.8), (3.4), and (3.5), respectively. Then
-(4.12a) $\lambda_{\min}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=0,$
-(4.12b) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=1-\delta_{\rm TG},$
+(4.1) $\lambda_{\max}\big(\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big)=1-\delta_{\rm TG},$
 where
@@
 Proof.
-The identity (4.12a) is obvious, because $\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\succeq 0$ is singular.
-Next, we focus on the proof of (4.12b).
-Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]$, it follows from (3.10) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$.
+Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)\subset[0,1]$, we get from (3.8) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$.
 By (3.7) and (3.8), we have
@@
 (4.4) $\operatorname*{rank}(X_{1})=\operatorname*{rank}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big).$
-We then prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
+Next, we prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
 (4.5) $\mathcal{N}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)\cap\mathcal{R}\big(A^{\frac{1}{2}}P\big)=\{0\}.$
-  • Recall that
-    $\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\quad\text{and}\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
-    If (4.16) holds, then
+  • Recall that $\mathcal{R}(\Pi)=\mathcal{R}\big(A^{\frac{1}{2}}P\big)$ and $\mathcal{N}(\Pi)=\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big)$. If (4.5) holds, then
     $\mathcal{N}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)\subseteq\mathcal{N}(\Pi),$
@@
     $\operatorname*{rank}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)=\operatorname*{rank}\big(P^{T}A^{\frac{1}{2}}\big)=\operatorname*{rank}(A_{\rm c})=s,$
-    which, together with (4.15), leads to $\operatorname*{rank}(X_{1})=s$.
+    which, together with (4.4), yields $\operatorname*{rank}(X_{1})=s$.
   • Conversely, if $\operatorname*{rank}(X_{1})=s$, then (4.5) holds. Otherwise, there must exist a vector $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(\Pi)$ such that
@@
     $\mathcal{N}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)=\mathcal{N}(\Pi),$
-    which follows from the relations $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$ and
+    which follows from the relations
+    $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)$
+    and
     $\operatorname*{rank}(\Pi)=s=\operatorname*{rank}(X_{1})=\operatorname*{rank}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big).$
@@
 (4.7) $\lambda\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)=\lambda\big(\widetilde{M}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big)=\lambda(\widetilde{M}A\Pi_{A}).$
-Combining (4.14), (4.17), and (4.18), we can arrive at the identity (4.12b). ∎
+Combining (4.3), (4.6), and (4.7), we can arrive at the identity (4.1). ∎
 Remark 4.2.
-The relation (4.16) will hold if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, since
-$\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A)\quad\text{and}\quad\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\subseteq\mathcal{R}(A).$
-Therefore, a sufficient condition for (4.16) to hold is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.6.
-We are now ready to present a two-sided estimate for $\|E_{\rm ITG}\|_{A}$.
+If $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, then (4.5) holds, since
+$\mathcal{N}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)=\mathcal{N}\big(\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big)=\mathcal{N}\big(\widetilde{M}A^{\frac{1}{2}}\big)=\mathcal{N}(A)$
+and
+$\mathcal{R}\big(A^{\frac{1}{2}}P\big)\subseteq\mathcal{R}(A)=\mathcal{N}(A)^{\perp}.$
+Obviously, a sufficient condition for (4.5) to be valid is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.7.
+We are now ready to present a convergence estimate for Algorithm 2.
 Theorem 4.3.
-Under the assumptions of Algorithm 2 and Lemma 4.2, the inexact two-grid convergence factor $\|E_{\rm ITG}\|_{A}$ satisfies
-(4.19) $\mathscr{L}(\beta_{2})\leq\|E_{\rm ITG}\|_{A}\leq\mathscr{U}(\beta_{1})$
-with
-$\mathscr{L}(\beta_{2})=\sqrt{1-\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}}$
+Let $\rho,\varepsilon\in[0,1)$ be two parameters. If
+(4.8) $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\rho\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}$
 and
-(4.20) $\mathscr{U}(\beta_{1})=\sqrt{1-\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}},$
-where $\widetilde{M}$, $\sigma_{\rm TG}$, $\beta_{1}$, $\beta_{2}$, and $\delta_{\rm TG}$ are given by (3.5), (3.8), (4.6a), (4.6b), and (4.13), respectively.
+(4.9) $\|\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}},$
+then
+(4.10) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}\big(1-\max\big\{\rho^{2}+\lambda_{n-r+1}(\widetilde{M}A),\,\delta_{\rm TG}\big\}\big)}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
+where $\widetilde{M}$ and $\delta_{\rm TG}$ are given by (3.4) and (4.2), respectively.
 Proof.
-In view of (4.10), one can get a two-sided estimate for $\|E_{\rm ITG}\|_{A}$ by bounding the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$.
-A lower bound for $\lambda_{n-r+1}(F_{1})$: By (4.11a), we have
-$\lambda_{n-r+1}(F_{1})$ $\geq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$=\lambda_{n-r+1}(\widetilde{M}A),$
-where we have used the facts (4.12a) and $\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}$. In addition, we have
-$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}(1-\beta_{1})A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{1}F_{\rm TG}\big{)}$
-$\geq(1-\beta_{1})\lambda_{1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{n-r+1}(F_{\rm TG})$
-$=\beta_{1}\sigma_{\rm TG}$
-and
-$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{1})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$\geq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{1})\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$=\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG}),$
-where, in the last equality, we have used the fact (4.12b). Hence,
-(4.21) $\lambda_{n-r+1}(F_{1})\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}.$
-An upper bound for $\lambda_{n-r+1}(F_{2})$: Using (4.11b) and (4.12b), we obtain
-$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{2}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$\leq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{2}\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$=\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}).$
-Moreover,
-$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$\leq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{2})\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$=\sigma_{\rm TG},$
-where, in the last equality, we have used the fact (4.12a). Thus,
-(4.22) $\lambda_{n-r+1}(F_{2})\leq\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}.$
-The estimate (4.19) then follows by combining (4.10), (4.21), and (4.22). ∎
+The condition (4.9) reads
+(4.11) $2\mathbf{r}_{\rm c}^{T}\hat{\mathbf{e}}_{\rm c}\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}.$
+From Algorithm 2, we have
+$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c}.$
+Then
+$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=(\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c})^{T}A(\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c})$
+$=\|\mathbf{u}-\mathbf{u}^{(1)}\|_{A}^{2}-2(\mathbf{u}-\mathbf{u}^{(1)})^{T}AP\hat{\mathbf{e}}_{\rm c}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}$
+$=\|\mathbf{u}-\mathbf{u}^{(1)}\|_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\hat{\mathbf{e}}_{\rm c}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}$
+[…]
+$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A(\mathbf{u}-\mathbf{u}^{(1)})-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
+$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A(\mathbf{u}-\mathbf{u}^{(1)})-(1-\varepsilon^{2})(\mathbf{u}-\mathbf{u}^{(1)})^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}(\mathbf{u}-\mathbf{u}^{(1)})$
+$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A^{\frac{1}{2}}\big(I-(1-\varepsilon^{2})\Pi\big)A^{\frac{1}{2}}(\mathbf{u}-\mathbf{u}^{(1)}).$
+Due to
+$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)(\mathbf{u}-\mathbf{u}^{(0)}),$
+it follows that
+(4.12) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq(\mathbf{u}-\mathbf{u}^{(0)})^{T}\big(E_{1}-(1-\varepsilon^{2})E_{2}\big)(\mathbf{u}-\mathbf{u}^{(0)}),$
+where
+$E_{1}=(I-MA)^{T}A(I-MA)\quad\text{and}\quad E_{2}=(I-MA)^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}(I-MA).$
+Observe that
+$E_{1}\succeq 0,\quad E_{2}\succeq 0,\quad\text{and}\quad E_{1}-E_{2}\succeq 0.$
+We next determine upper bounds for $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2}$, $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2}$, and their difference.
+  • Since
+    $E_{1}=A^{\frac{1}{2}}\big(I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big)A^{\frac{1}{2}},$
+    we obtain
+[…]
+    A straightforward analysis yields
+    (4.13) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2}\leq\big(1-\lambda_{n-r+1}(\widetilde{M}A)\big)\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2},$
+    where we have used the fact
+    $\lambda(\overline{M}A)=\lambda(\widetilde{M}A).$
+  • It is easy to see that
+    $E_{2}=A^{\frac{1}{2}}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)A^{\frac{1}{2}}.$
+    In light of (4.1), we have
+    (4.14) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2}\leq(1-\delta_{\rm TG})\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2}.$
+  • Note that
+    $E_{1}-E_{2}$ $=(I-MA)^{T}A^{\frac{1}{2}}(I-\Pi)A^{\frac{1}{2}}(I-MA)$
+    $=(I-MA)^{T}A^{\frac{1}{2}}(I-\Pi)^{T}(I-\Pi)A^{\frac{1}{2}}(I-MA)$
+    $=(I-MA)^{T}(I-\Pi_{A})^{T}A(I-\Pi_{A})(I-MA)$
+    $=E_{\rm TG}^{T}AE_{\rm TG}.$
+    By (2.7) and (4.8), we have
+    (4.15) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}=\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}^{2}\leq\rho^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2}.$
+According to (4.12), we deduce that
+$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}+\varepsilon^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2},$
+which, combined with (4.14) and (4.15), leads to
+(4.16) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}(1-\delta_{\rm TG})}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}.$
+In addition, we get from (4.12) that
+$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq(1-\varepsilon^{2})\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}+\varepsilon^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2},$
+which, together with (4.13) and (4.15), yields
+(4.17) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}\big(1-\rho^{2}-\lambda_{n-r+1}(\widetilde{M}A)\big)}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}.$
+The desired estimate then follows immediately by combining (4.16) and (4.17). ∎
 Remark 4.4.
-By (4.11a) and (4.12b), we have
-$1-\delta_{\rm TG}$ $=\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
-$=\lambda_{n}\big{(}F_{\rm TG}-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
-$\geq\lambda_{n-r+1}(F_{\rm TG})-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
-$=\sigma_{\rm TG}-\lambda_{n-r+1}(\widetilde{M}A),$
-which gives
-$\sigma_{\rm TG}\leq 1-\delta_{\rm TG}+\lambda_{n-r+1}(\widetilde{M}A).$
-According to (4.6a) and (4.6b), we deduce that $0<\beta_{1}\leq\beta_{2}\leq 1$.
-  • (i)
-    If $\beta_{2}\rightarrow 0^{+}$, then $F_{i}\rightarrow A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}$ for $i=1,2$, which, together with (4.10), yields
-    $\|E_{\rm ITG}\|_{A}\rightarrow\sqrt{1-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}}=\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
-    This asymptotic behavior can also be observed from (4.19) as $\beta_{2}\rightarrow 0^{+}$.
-  • (ii)
-    In the special case $B_{\rm c}=A_{\rm c}$ (i.e., $\beta_{1}=\beta_{2}=1$), we have
-    $\mathscr{L}(1)=\mathscr{U}(1)=\sqrt{1-\sigma_{\rm TG}},$
-    which leads to the convergence identity (3.7).
+Recall that
+$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
+where $\|E_{\rm TG}\|_{A}<1$.
+For the sake of practicality, we have made a more general assumption in Theorem 4.3.
+Another main assumption made in Theorem 4.3 is (4.9), which, in fact, characterizes the relative accuracy of the coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$.
 Remark 4.5.
@@
 which will be satisfied if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$.
-4.2. The general case
-In this subsection, we present a convergence estimate for Algorithm 2 with general coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$, as described in the following theorem.
-Theorem 4.10.
-Under the assumptions of Algorithm 2, if
-(4.23) $\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}$
-for some $\varepsilon\in[0,1)$, then
-(4.24) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\mathscr{U}(1-\varepsilon^{2})\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
-where $\mathscr{U}(\cdot)$ is given by (4.20).
-Proof.
-The condition (4.23) reads
-$2\big{(}A_{\rm c}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}\big{)}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
-Note that $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
-$\mathbf{r}_{\rm c}=P^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}\in\mathcal{R}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\mathcal{R}(A_{\rm c}).$
-We then have
-(4.25) $2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
-From Algorithm 2, we have
-$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket.$
-Then
-$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}$
-$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}AP\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
-$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
-$\leq\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}$
-$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
-$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}$
-$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\big{(}I-(1-\varepsilon^{2})\Pi\big{)}A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)},$
-where we have used the inequality (4.25). Due to
-$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
-and
-$\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}},$
-it follows that
-$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}^{T}A^{\frac{1}{2}}(I-F)A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
-with
-$F=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+(1-\varepsilon^{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
-Then
-$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\bigg{(}1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}\bigg{)}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}^{2}.$
-Similarly to the proof of Lemma 3.1, one can prove that
-$\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}=\lambda_{n-r+1}(F).$
-Thus,
-$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F)}\,\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}.$
-The estimate (4.24) can be obtained by bounding $\lambda_{n-r+1}(F)$ from below. Indeed, it holds that
-$\lambda_{n-r+1}(F)\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,(1-\varepsilon^{2})\sigma_{\rm TG},\,\sigma_{\rm TG}-\varepsilon^{2}(1-\delta_{\rm TG})\big{\}};$
-see the proof of (4.21). This completes the proof. ∎
 Remark 4.6.
 From Algorithms 1 and 2, we have
-$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\big{\|}P\big{(}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}\big{\|}_{A}=\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}},$
-i.e., the perturbation bound of two-grid solution is equal to that of coarse solution. Therefore, the accuracy of two-grid solver can be controlled by that of coarse solver.
+$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\|P(\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c})\|_{A}=\|\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}},$
+i.e., the perturbation bound of two-grid solution is equal to that of coarse solution. Thus, the accuracy of two-grid solver can be controlled by that of coarse solver.
+Let $\rho_{\rm TG}$ and $\rho_{\rm ITG}$ denote the asymptotic convergence factors of exact and inexact two-grid methods, respectively. The estimate (4.10) yields
+$\rho_{\rm ITG}\leq\sqrt{\rho_{\rm TG}^{2}+\varepsilon^{2}\big(1-\max\big\{\rho_{\rm TG}^{2}+\lambda_{n-r+1}(\widetilde{M}A),\,\delta_{\rm TG}\big\}\big)},$
+which implies that
+(4.18) $\rho_{\rm ITG}\leq\sqrt{\rho_{\rm TG}^{2}+\varepsilon^{2}\big(1-\rho_{\rm TG}^{2}\big)}.$
+If $\rho_{\rm TG}$ can be bounded uniformly from above ($\rho_{\rm TG}\leq C$ for some constant $C<1$), then
+$\rho_{\rm ITG}\leq\sqrt{\varepsilon^{2}+(1-\varepsilon^{2})\rho_{\rm TG}^{2}}\leq\sqrt{\varepsilon^{2}+(1-\varepsilon^{2})C^{2}}<1\quad\forall\,\varepsilon\in[0,1),$
+that is, for any fixed $\varepsilon\in[0,1)$, $\rho_{\rm ITG}$ can also be bounded uniformly.
+In addition, we deduce from (4.18) that
+$\rho_{\rm ITG}\leq\rho_{\rm TG}\sqrt{1+\frac{\varepsilon^{2}\big(1-\rho_{\rm TG}^{2}\big)}{\rho_{\rm TG}^{2}}}.$
+This may serve as a guide for designing some practical algorithms. For example, if $\rho_{\rm TG}=0.3$, then $\rho_{\rm ITG}\leq 0.4$ can be achieved by setting $\varepsilon=\frac{1}{\sqrt{13}}$.
+4.2. Numerical experiments
+In order to intuitively illustrate the estimate (4.18), we next provide a numerical example.
+Consider solving the 2D Poisson’s equation with homogeneous Neumann boundary conditions on the unit square $\Omega=(0,1)\times(0,1)$:
+(4.19) $\left\{\begin{aligned} -\Delta u&=f\quad\text{in $\Omega$},\\ \frac{\partial u}{\partial n}&=0\quad\text{on $\partial\Omega$}.\end{aligned}\right$.
+The source term $f=f(x,y)$ should satisfy the following compatibility condition:
+$\int_{\Omega}f(x,y)\,\mathrm{d}x\mathrm{d}y=0.$
+To solve the problem (4.19) numerically, we partition the domain $\Omega$ uniformly and set
+$(x_{i},y_{j})=(ih,jh)\quad\forall\,i,j=0,1,\ldots,m+1,$
+where $h=1/(m+1)$. A common way to discretize the problem (4.19) involves the following ghost points:
+$(x_{-1},y_{j})$ $=(-h,jh)\quad j=0,\ldots,m+1;$
+$(x_{m+2},y_{j})$ $=(1+h,jh)\quad j=0,\ldots,m+1;$
+$(x_{i},y_{-1})$ $=(ih,-h)\quad i=0,\ldots,m+1;$
+$(x_{i},y_{m+2})$ $=(ih,1+h)\quad i=0,\ldots,m+1;$
+see the yellow points in Figure 1.
+Figure 1. The restriction and prolongation processes between fine and coarse grids
+The purpose of introducing these ghost points is to produce approximations at the boundary points.
+With this extended grid, we can apply the standard five-point difference scheme to the Laplacian operator at both interior and boundary points.
+Moreover, these ghost points will be used to form central difference approximations to the normal derivatives.
+These choices lead to the following discrete equations:
+$\left\{\begin{aligned} 4u_{i,j}-u_{i-1,j}-u_{i+1,j}-u_{i,j-1}-u_{i,j+1}&=h^{2}f_{i,j}\quad i,j=0,\ldots,m+1,\\ u_{1,j}-u_{-1,j}&=0\quad j=0,\ldots,m+1,\\ u_{m+2,j}-u_{m,j}&=0\quad j=0,\ldots,m+1,\\ u_{i,1}-u_{i,-1}&=0\quad i=0,\ldots,m+1,\\ u_{i,m+2}-u_{i,m}&=0\quad i=0,\ldots,m+1,\end{aligned}\right$.
+where $u_{i,j}\approx u(x_{i},y_{j})$ and $f_{i,j}=f(x_{i},y_{j})$. Eliminating the ghost unknowns $u_{-1,j}$, $u_{m+2,j}$, $u_{i,-1}$, and $u_{i,m+2}$ yields
+$\left\{\begin{aligned} 8u_{i,j}-2u_{i-1,j}-2u_{i+1,j}-2u_{i,j-1}-2u_{i,j+1}&=2h^{2}f_{i,j}\quad i,j=1,\ldots,m,\\ 4u_{i,0}-u_{i-1,0}-u_{i+1,0}-2u_{i,1}&=h^{2}f_{i,0}\quad i=1,\ldots,m,\\ 4u_{0,j}-2u_{1,j}-u_{0,j-1}-u_{0,j+1}&=h^{2}f_{0,j}\quad j=1,\ldots,m,\\ 4u_{m+1,j}-2u_{m,j}-u_{m+1,j-1}-u_{m+1,j+1}&=h^{2}f_{m+1,j}\quad j=1,\ldots,m,\\ 4u_{i,m+1}-u_{i-1,m+1}-u_{i+1,m+1}-2u_{i,m}&=h^{2}f_{i,m+1}\quad i=1,\ldots,m,\\ 2u_{0,0}-u_{1,0}-u_{0,1}&=\frac{1}{2}h^{2}f_{0,0},\\ 2u_{m+1,0}-u_{m,0}-u_{m+1,1}&=\frac{1}{2}h^{2}f_{m+1,0},\\ 2u_{0,m+1}-u_{1,m+1}-u_{0,m}&=\frac{1}{2}h^{2}f_{0,m+1},\\ 2u_{m+1,m+1}-u_{m,m+1}-u_{m+1,m}&=\frac{1}{2}h^{2}f_{m+1,m+1}.\end{aligned}\right$.
+Let
+$B$ $=\operatorname*{diag}(1,2,\ldots,2,1)\in\mathbb{R}^{(m+2)\times(m+2)},$
+$C$ $={\rm tridiag}(-1,4,-1)-2\big(\mathbf{e}_{1}\mathbf{e}_{1}^{T}+\mathbf{e}_{m+2}\mathbf{e}_{m+2}^{T}\big)\in\mathbb{R}^{(m+2)\times(m+2)},$
+$D$ $={\rm tridiag}(-1,0,-1)\in\mathbb{R}^{(m+2)\times(m+2)},$
+$\bm{\omega}$ $=(1,2,\ldots,2,1)^{T}\in\mathbb{R}^{m+2},$
+$\mathbf{f}_{0}$ $=h^{2}\big(f_{0,0},f_{1,0},\ldots,f_{m+1,0},\ldots,f_{0,m+1},f_{1,m+1},\ldots,f_{m+1,m+1}\big)^{T}\in\mathbb{R}^{(m+2)^{2}},$
+where $\mathbf{e}_{1}$ and $\mathbf{e}_{m+2}$ denote the first and last columns of the identity matrix $I_{m+2}$, respectively.
+By ordering all $u_{i,j}$ lexicographically, we end up with the linear system
+$A\mathbf{u}=\mathbf{f},$
+where
+$A=B\otimes C+D\otimes B\quad\text{and}\quad\mathbf{f}=\frac{1}{2}(\bm{\omega}\otimes\bm{\omega})\circ\mathbf{f}_{0}.$
+The symbols $\otimes$ and $\circ$ denote the Kronecker and Hadamard products of two matrices, respectively.
+Assume that $m\geq 1$ is an odd number.
+We choose the points $(x_{i},y_{j})$ with indices $(i,j)\in\{0,2,\ldots,m+1\}\times\{0,2,\ldots,m+1\}$ as coarse ones; see the black points in Figure 1.
+Main settings in our numerical experiments are listed as follows:
+  • $f$ is simply chosen as zero;
+  • $\mathbf{u}^{(0)}$ is chosen randomly and then fixed;
+  • $M$ is chosen as the symmetrized Gauss–Seidel type;
+  • $R=R_{0}\otimes R_{0}$, where
+    $R_{0}=\frac{1}{4}\begin{pmatrix}2&1&&&&&&&\\ &1&2&1&&&&&\\ &&&&\ddots&&&&\\ &&&&&1&2&1&\\ &&&&&&&1&2\end{pmatrix}\in\mathbb{R}^{\frac{m+3}{2}\times(m+2)};$
+  • $P=4R^{T}$, which is based on the prolongation stencil shown in Figure 1;
+  • $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$ is chosen as the standard conjugate gradient method.
+The corresponding numerical results are listed in Table 1 below.
+$h$ $1/32$ $1/64$ $1/128$ $1/256$ $1/512$
+$\rho_{\rm TG}$ $0.2231$ $0.2238$ $0.2236$ $0.2238$ $0.2238$
+$\varepsilon=0.1$ $\rho_{\rm ITG}$ $0.2231$ $0.2238$ $0.2236$ $0.2238$ $0.2238$
+Upper bound in (4.18) $0.2435$ $0.2441$ $0.2439$ $0.2441$ $0.2441$
+$\varepsilon=0.2$ $\rho_{\rm ITG}$ $0.2231$ $0.2238$ $0.2236$ $0.2238$ $0.2238$
+Upper bound in (4.18) $0.2963$ $0.2968$ $0.2966$ $0.2968$ $0.2968$
+$\varepsilon=0.3$ $\rho_{\rm ITG}$ $0.2598$ $0.2848$ $0.2913$ $0.2947$ $0.2963$
+Upper bound in (4.18) $0.3678$ $0.3682$ $0.3681$ $0.3682$ $0.3682$
+$\varepsilon=0.4$ $\rho_{\rm ITG}$ $0.3285$ $0.3816$ $0.3852$ $0.3915$ $0.3946$
+Upper bound in (4.18) $0.4492$ $0.4495$ $0.4494$ $0.4495$ $0.4495$
+$\varepsilon=0.5$ $\rho_{\rm ITG}$ $0.4500$ $0.4669$ $0.4833$ $0.4918$ $0.4932$
+Upper bound in (4.18) $0.5360$ $0.5363$ $0.5362$ $0.5363$ $0.5363$
+Table 1. Asymptotic convergence factors and their upper bounds
+From Table 1, one can observe that
+  • $\rho_{\rm TG}$ is essentially independent of grid size, and $\rho_{\rm ITG}$ can be bounded uniformly for any fixed $\varepsilon\in[0,1)$;
+  • Satisfactory convergence may be achieved even if low-precision coarse solvers are used (so it may not be necessary to solve the coarse problem very precisely in practice);
+  • Our theory can provide a good prediction for the actual convergence factors.
 5. Conclusions
-Two-grid convergence theory plays an important role in the design and analysis of multigrid methods.
-In this paper, we have established a new convergence theory for two-grid methods applied to SPSD linear systems.
-In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for the convergence factor of two-grid methods.
-More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers (both linear and general types are considered).
-A distinctive feature in comparison with the existing ones is that our theory does not require any additional assumptions on the system matrix, especially on its null space.
+Two-grid theory plays a fundamental role in the design and analysis of multigrid methods.
+In this paper, we establish a new two-grid convergence theory for SPSD systems, which does not require any additional conditions on the coefficient matrix.
+When the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for characterizing the convergence factor of two-grid methods.
+More generally, we present a convergence estimate for two-grid methods with approximate coarse solvers.
+A numerical example is also provided to illustrate our theoretical estimate.
 Acknowledgments
@@
 12401479), the Natural Science Foundation of Jiangsu Province (Grant No.
-BK20241257), and the Scientific Research Foundation of the Southeast University (Grant No.
-4007012324).
+BK20241257), and the Start-up Research Fund of Southeast University (Grant No.
+RF1028623372).